### v1

# Qronos: Correcting the Past by Shaping the Future… in Post-Training Quantization

Abstract
We introduce Qronos—a new state-of-the-art post-training quantization algorithm that sequentially rounds and updates neural network weights. Qronos not only explicitly corrects errors due to both weight and activation quantization, but also errors resulting from quantizing previous layers. Our iterative algorithm is based on an interpretable and disciplined optimization framework that subsumes and surpasses existing data-driven approaches. At each step, Qronos alternates between error correction and diffusion via optimal update rules. Importantly, we prove that Qronos admits an efficient implementation that uses the Cholesky decomposition for solving least-squares problems. We also demonstrate that Qronos is compatible with existing transformation techniques such as Hadamard-based incoherence processing and weight-activation scaling equalization, among others. We evaluate Qronos using recent autoregressive language generation models in the Llama3 family; Qronos consistently outperforms previous state-of-the-art adaptive rounding methods when quantizing the weights, activations, and/or KV caches.
1 Introduction
Recent advances in post-training quantization (PTQ) have enabled the practical use of few-bit weights and activations for large language model (LLM) inference, typically by focusing on one or both aspects of the quantization pipeline, visualized in Figure 1. The first aspect involves modifying the weights and activations of a model to make them more amenable to quantization, often through transformations that exploit invariances within the compute graph. The second aspect more directly concerns the design of the quantization mapping itself; it involves using data to minimize quantization error by either calibrating the quantization grid—defined by a bit width, scaling factor, and zero point—or adaptively rounding the (potentially transformed) weights.
The latest innovations in PTQ, including (Ashkboos et al., 2024; Liu et al., 2025), among many others, are skewed towards proposing and improving transformations that address the quantization challenges exacerbated in LLMs. These studies often only consider round-to-nearest (RTN) and OPTQ (Frantar et al., 2023), also known as GPTQ. Meanwhile, our work explicitly focuses on improving the rounding method while remaining compatible with these transformations to collectively improve the quantization pipeline. We are in part motivated by the latest data-driven rounding algorithms, namely OPTQ (Frantar et al., 2023) and GPFQ (Lybrand and Saab, 2021). OPTQ, which is widely used in the literature, is a sequential algorithm that alternates between rounding the current weight and adjusting future unquantized weights to reduce error, a process we refer to as diffusion. However, OPTQ does not correct for quantization error in the activations or for error introduced by quantizing previous layers. In contrast, GPFQ explicitly corrects for quantization error in both the weights and activations in each rounding iteration, as well as error from previous layers, but does not adjust the future unquantized weights.
Contributions. We introduce Qronos—a new state-of-the-art post-training quantization algorithm that sequentially rounds and updates neural network weights. Qronos demonstrably subsumes and surpasses OPTQ and GPFQ, explicitly correcting quantization error in both the weights and activations of previous layers while diffusing error into future weights. In Section 3, we start by presenting Qronos in a well-disciplined and mathematically interpretable form, then rigorously derive an equivalent computationally efficient implementation. Our implementation exploits existing optimizations for OPTQ and GPFQ, such as Cholesky decomposition and block-level error diffusion. Indeed, in Section 4.3, we show our reductions significantly improve algorithm scaling, yielding up to a 13.8$\times$ improvement in algorithm runtime for a single-layer microbenchmark. As a theoretical by-product, we also derive a novel interpretation of OPTQ in Appendix D, which shows that it greedily corrects the cumulative quantization error incurred over previous iterations.
We empirically evaluate Qronos using the Llama3 model family (Grattafiori et al., 2024) and compare against RTN, OPTQ, and GPFQ while demonstrating compatibility with notable transformations for both weight-only quantization (e.g., SmoothQuant (Xiao et al., 2023) and incoherence processing Chee et al. (2023)) and weight-activation quantization (e.g., QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2025)). To our knowledge, our work is the first to isolate the impact of the rounding algorithm using a fixed quantization grid for each transformation method (or lack thereof).
Our experiments show that Qronos consistently yields marked improvement over existing rounding methods; for example, when quantizing the weights of Llama3 models to 2 bits or fewer, as shown in Table 1, Qronos provides massive improvements in perplexity over RTN, OPTQ, and GPFQ.
Table 1: Weight-only quantization of Llama3 models. We jointly apply Hadamard-based incoherence processing (Ashkboos et al., 2024) and weight magnitude reduction (Zhang et al., 2024) as quantization transforms (stage 1) and compare different rounding functions (stage 2).
WikiText2 ($\downarrow$) 0-shot ($\uparrow$)
1B 3B 8B     1B 3B 8B
BF16 - 8.9 7.1 5.9     59.4 67.5 74.4
2-bit RTN 1e4 1e4 1e4     32.2 32.1 32.9
OPTQ 42.5 20.4 19.0     35.4 40.6 40.7
GPFQ 48.3 22.8 18.0     34.4 38.3 39.8
Qronos 24.6 14.9 12.4     37.8 42.0 45.3
1.58-bit RTN 2e5 3e5 6e5     32.0 32.0 31.9
OPTQ 4e3 6e2 3e2     32.0 32.2 32.4
GPFQ 7e2 8e2 6e2     32.4 32.6 32.3
Qronos 79.5 48.3 34.8     32.7 32.9 33.5
2 Background and Related Work
Figure 1: The modern quantization pipeline is typically a two-stage process consisting of (1) transformations that make weights and/or activations more amenable to quantization, followed by (2) rounding functions that map weights and/or activations onto a quantization grid.
Before presenting the details of our approach, we first provide a short review of prior works that focus on the two key aspects of quantization we have mentioned: transformation techniques and rounding schemes. Figure 1 illustrates how these two categories interact within the quantization pipeline.
Methods based on transformations. Many recent works propose transformations of weights and/or activations to facilitate quantization. One line of work, initially proposed for MobileNets (Nagel et al., 2019), exploits scaling invariance in neural network compute graphs to equalize the range or precision of weights and activations before quantization. Recent variants leverage scale invariance to redistribute quantization difficulty between weights and activations, with various proposals for learning scales or ranges based on custom objective functions (Xiao et al., 2023; Shao et al., 2024; Lin et al., 2024). Another line of work uses rotations within a compute graph to normalize weight and activation distributions, initially leveraging random orthogonal rotations to promote weight incoherence (Chee et al., 2023). Recent variants employ efficient Hadamard rotations (Tseng et al., 2024a; Ashkboos et al., 2024), Stiefel manifold optimizations (Liu et al., 2025; Hu et al., 2025), and rotation expansion techniques (Adepu et al., 2024; Franco et al., 2025a). Finally, distinct from these invariance-based approaches, MagR (Zhang et al., 2024) directly minimizes the $\ell_{\infty}$ norm of weights via proximal gradient descent to reduce dynamic range before quantization. While we do not introduce novel transformations of this type in this work, we demonstrate that existing transformations can be combined with our proposed method.
Methods based on rounding. The earliest line of work on rounding relies on continuous optimization strategies based on gradient descent (Nagel et al., 2020). Although more recent methods exist (Hubara et al., 2021; Li et al., 2021), they are not commonly evaluated on LLMs due to computational cost. Early work on LLMs focused on grid scaling or shifting to reduce weight quantization error; for example, LLM.int8() (Dettmers et al., 2022) and ZeroQuant (Yao et al., 2022) directly round to nearest after heuristically selecting the quantization grid, i.e., bit width, scaling factors and zero points. Recent variants have explored more intricate quantization grids, e.g., vector quantization (Tseng et al., 2024a) or trellis codes (Tseng et al., 2024b), that exploit knowledge or assumptions of the data. The most relevant line of work to ours adopts principled discrete optimization using greedy, sequential rounding strategies to select quantized weights to minimize the layer-wise reconstruction error, and includes OBQ (Frantar and Alistarh, 2022), OPTQ (Frantar et al., 2023), GPFQ (Lybrand and Saab, 2021; Zhang et al., 2023), and LDLQ (Chee et al., 2023). Qronos falls within this category.
Notation. Throughout the paper, $\mathcal{A}$ denotes the discrete quantization grid (or alphabet) used for weight quantization, and $\mathcal{Q}$ denotes the corresponding RTN operator associated with $\mathcal{A}$. The weight matrix of a layer is denoted by $W\in\mathbb{R}^{N\times N^{\prime}}$, where each of the $N^{\prime}$ columns represents a $N$-dimensional channel. Depending on the setting, different channels (i.e., columns) of the weight matrix may have different quantization grids and RTN operators due to scaling factor granularity (e.g., per-channel or per-group quantization). With a slight abuse of notation, we use the same symbols $\mathcal{A}$ and $\mathcal{Q}$ for all entries. When quantizing $W$, we use $X\in\mathbb{R}^{m\times N}$ to denote the input calibration dataset of $m$ samples (e.g., tokens) for the layer, resulting from the original pre-trained model, and $\widetilde{X}\in\mathbb{R}^{m\times N}$ to denote the input calibration dataset coming from the partially quantized model.
Given a vector $v\in\mathbb{R}^{n}$, we use $v_{i}$ for its $i$-th entry, $v_{\geq j}$ for the subvector $(v_{j},\dots,v_{n})^{\top}$, and we define $v_{\leq j}$ analogously. $\|v\|$ is the Euclidean norm of $v$. Given a matrix $A\in\mathbb{R}^{m\times n}$, we use $A_{i}$ to denote its $i$-th column. We use $A_{\geq j}$ to denote the submatrix $(A_{j},\dots,A_{n})$. Similarly, we use $A_{\mathcal{I},\mathcal{J}}$ to represent the submatrix of $A$ consisting of rows indexed by $\mathcal{I}$ and columns indexed by $\mathcal{J}$. For example, $A_{\geq 2,\geq 2}$ will denote the submatrix of $A$ obtained by removing the first row and the first column. $\|A\|_{F}$ denotes the Frobenius norm of $A$. We use $\text{col}(A)$ to denote the column space of $A$. $P_{A}$ is the orthogonal projection onto $\text{col}(A)$, and $P_{A^{\perp}}$ the projection onto its orthogonal complement. Throughout this paper, all indices start from 1, following the standard mathematical convention.
Layer-wise reconstruction and error correction. Data-driven weight quantization methods typically aim to approximately minimize [^1] Equation 1 is an instance of integer least-squares problems, which are known to be NP-hard (Hassibi and Vikalo, 2002). Thus, the best that one can hope for are approximate solutions. the layer-wise reconstruction error given by
$\min_{Q\in\mathcal{A}^{N\times N^{\prime}}}\|XW-XQ\|_{F}^{2}.$ (1)
At an arbitrary layer the goal is to compute a quantized weight matrix $Q\in\mathcal{A}^{N\times N^{\prime}}$ by approximately solving (1) to preserve the output activations $XW$ under quantization. In practice, however, quantizing weights in earlier layers affects the input to subsequent layers. Let $\widetilde{X}\in\mathbb{R}^{m\times N}$ denote the activation matrix produced by a partially quantized model, where earlier layers have already been quantized. To account for the propagation of quantization error, we use a modified formulation, instead of (1), that targets the mismatch between the original output $XW$ and $\widetilde{X}Q$ by approximately solving $\min_{Q\in\mathcal{A}^{N\times N^{\prime}}}\|XW-\widetilde{X}Q\|_{F}^{2}.$ The type of mismatch in this formulation is typically not addressed in the literature but arises naturally in both weight-only and weight-activation quantization settings. For instance, in weight-only quantization, $\widetilde{X}$ arises as the output of previously quantized layers, while in weight-activation quantization, one may encounter $\mathcal{Q}(\widetilde{X})$ rather than $\widetilde{X}$ if activations are quantized. Throughout this paper, we use the notation $(X,\widetilde{X})$ to refer generically to mismatched input pairs.
3 Qronos
We begin by describing the iterations associated with Qronos in Section 3.1. The iterations follow a disciplined and mathematically interpretable framework that alternates between error correction and diffusion using optimal update rules. We then prove that the explicit solutions to these minimization problems admit the efficient implementation detailed in Algorithm 1. In Section 3.2, we provide deeper intuition behind Qronos in the context of previous state-of-the-art rounding algorithms, namely GPFQ and OPTQ. Meanwhile, we also derive a novel interpretation of OPTQ in Appendix D, which shows that it corrects the cumulative quantization error incurred over all the previous iterations.
Algorithm 1 Qronos
$H^{-1}=(\widetilde{X}^{\top}\widetilde{X})^{-1}=LL^{\top}$
for every $w$ in $W$ (in parallel) do
   $q=\mathbf{0}^{N}$
   $w^{(0)}\leftarrow\text{copy}(w)$
   $q_{1}=\mathcal{Q}\left(\dfrac{\widetilde{X}_{1}^{\top}(Xw-\widetilde{X}_{\geq 2}w^{(0)}_{\geq 2})}{\|\widetilde{X}_{1}\|^{2}}\right)$triangleright Calc. $q_{1}$ from (6), use (14) for a memory efficient ver.
   $w^{(1)}_{\geq 2}=\widetilde{X}_{\geq 2}^{\dagger}\left(Xw-\widetilde{X}_{1}q_{1}\right)$triangleright Calc. $w^{(1)}_{\geq 2}$ from (7), see (15) for a memory efficient ver.
   for $t=2$ to $N$ do $\triangleright$ Efficient implementation (Lemma 3.2)
     $q_{t}=\mathcal{Q}(w^{(t-1)}_{t})$
     $w^{(t)}_{\geq t+1}=w^{(t-1)}_{\geq t+1}-L_{\geq t+1,t}\cdot(w^{(t-1)}_{t}-q_{t})/L_{tt}$
   end for
end for
return $Q$
3.1 Algorithm and Efficient Implementation
Let us first note that Qronos can process each column $w\in\mathbb{R}^{N}$ of $W\in\mathbb{R}^{N\times N^{\prime}}$ independently and in parallel to produce each column $q\in\mathcal{A}^{N}$ of $Q\in\mathcal{A}^{N\times N^{\prime}}$. Ideally, the goal is to find $q$ that minimizes $\frac{1}{2}\|Xw-\widetilde{X}q\|^{2}$. Since this problem is NP-hard, we propose an efficient sequential algorithm to approximate its solution. At each iteration, Qronos first selects the quantized weight that optimally corrects the current approximation error, holding the remaining weights fixed; see (2) below. It then updates the unquantized weights to optimally compensate for the rounding error, a process we refer to as error diffusion; see (3).
Let $w$, without superscripts or subscripts, denote the original unquantized neuron. After determining $q_{t-1}$, let $w^{(t-1)}_{\geq t}$ represent the updated unquantized weights corresponding to indices $t$ through $N$. The full state of the algorithm after step $t-1$ is thus given by the vector $w^{(t-1)}=(q_{\leq t-1},w^{(t-1)}_{\geq t}),$ with the initialization $w^{(0)}=w$. At step $t$, the algorithm alternates between selecting $q_{t}$ through error correction and updating the remaining weights through error diffusion. The update rules are given by
$q_{t}$ $=\underset{p\in\mathcal{A}}{\mathrm{argmin}}\,\frac{1}{2}\|Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-p\widetilde{X}_{t}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j}\|^{2},$ (2)
$w^{(t)}_{\geq t+1}$ $=\underset{(v_{t+1},\dots,v_{N})\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\,\frac{1}{2}\|Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}v_{j}\widetilde{X}_{j}\|^{2}.$ (3)
These optimization problems admit the following closed-form solutions (see Proposition A.1):
$q_{t}$ $=\mathcal{Q}\left(\frac{\left\langle Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j},\widetilde{X}_{t}\right\rangle}{\|\widetilde{X}_{t}\|^{2}}\right),$ (4)
$w^{(t)}_{\geq t+1}$ $=\widetilde{X}_{\geq t+1}^{\dagger}\left(Xw-\widetilde{X}_{\leq t}q_{\leq t}\right).$ (5)
While these expressions follow directly from the optimization problems, computing $q_{t}$ and $w^{(t)}_{\geq t+1}$ in this form is not computationally efficient and scales poorly, as we will show in Section 4.3. To address this, we present Theorem 3.1, which shows that for all $t\geq 2$, $q_{t}$ can be computed via RTN, enabling a simpler implementation. In Lemma 3.2, we further show that the update for $w^{(t)}_{\geq t+1}$ also admits an efficient implementation using Cholesky decomposition to solve the associated least-squares problem. Together, these results yield the practical and scalable implementation summarized in Algorithm 1.
Theorem 3.1.
Let $(q_{t},w^{(t-1)}_{\geq t})$ be the iterates generated by (2) and (3), with initialization $w^{(0)}_{\geq 1}=w$. Define an alternative sequence $(\hat{q}_{t},\hat{w}^{(t-1)}_{\geq t})$ using the same initialization $\hat{w}^{(0)}_{\geq 1}=w$, by setting
$\hat{q}_{1}$ $=\underset{p\in\mathcal{A}}{\arg\min}\,\frac{1}{2}\|Xw-p\widetilde{X}_{1}-\sum_{j=2}^{N}w_{j}\widetilde{X}_{j}\|^{2},$ (6)
$\hat{w}^{(1)}_{\geq 2}$ $=\underset{(v_{2},\dots,v_{N})\in\mathbb{R}^{N-1}}{\arg\min}\,\frac{1}{2}\|Xw-\hat{q}_{1}\widetilde{X}_{1}-\sum_{j=2}^{N}v_{j}\widetilde{X}_{j}\|^{2},$ (7)
and, for $t=2,\dots,N$, define
$\hat{q}_{t}$ $=\mathcal{Q}(\hat{w}^{(t-1)}_{t}),$ (8)
$\hat{w}^{(t)}_{\geq t+1}$ $=\underset{(v_{t+1},\dots,v_{N})\in\mathbb{R}^{N-t}}{\arg\min}\,\frac{1}{2}\|(\hat{q}_{t}-\hat{w}^{(t-1)}_{t})\widetilde{X}_{t}+\sum_{j=t+1}^{N}(v_{j}-\hat{w}^{(t-1)}_{j})\widetilde{X}_{j}\|^{2}.$ (9)
Then for $t=1,\dots,N$, the two procedures yield identical iterates: $(q_{t},w^{(t-1)}_{\geq t})=(\hat{q}_{t},\hat{w}^{(t-1)}_{\geq t})$.
Starting from the second iteration, Theorem 3.1 shows that the updates in (2) and (3) can be equivalently reformulated as (8) and (9), respectively. This reformulation allows $q_{t}$ to be obtained via RTN for $t\geq 2$, followed by an adjustment of the remaining weights using a single activation matrix $\widetilde{X}$ to compensate for the one-step quantization error $(q_{t}-w^{(t-1)}_{t})\widetilde{X}_{t}$.
To further accelerate this adjustment step, we now present Lemma 3.2, which establishes the equivalence of the update in (9) (for $t\geq 2$) with a Cholesky-based least-squares solution [^2] We do not claim that Lemma 3.2 is novel, though we were unable to find it stated explicitly in the literature.. For notational simplicity, we slightly abuse the indexing by treating $t=2$ as a ‘restart.’
Lemma 3.2 (Equivalence of Least-Squares Formulation and Cholesky Formulation).
Assume that $H=X^{\top}X$ is invertible, and let $H^{-1}=LL^{\top}$ denote its Cholesky decomposition, with $L$ lower triangular. Then, starting from $w^{(0)}=w$, the update rules
$q_{t}$ $=\mathcal{Q}(w^{(t-1)}_{t}),$ (10)
$w^{(t)}_{\geq t+1}$ $=\underset{(v_{t+1},\dots,v_{N})\in\mathbb{R}^{N-t}}{\arg\min}\,\frac{1}{2}\|(q_{t}-w^{(t-1)}_{t})X_{t}+\sum_{j=t+1}^{N}(v_{j}-w^{(t-1)}_{j})X_{j}\|^{2}$ (11)
are equivalent to the Cholesky-based iterations
$q_{t}$ $=\mathcal{Q}(w^{(t-1)}_{t}),$ (12)
$w^{(t)}_{\geq t+1}$ $=w^{(t-1)}_{\geq t+1}+\Delta^{(t)},$ (13)
where
$\Delta^{(t)}=-(w^{(t-1)}_{t}-q_{t})\frac{L_{\geq t+1,\,t}}{L_{tt}}\in\mathbb{R}^{N-t}.$
Lemma 3.2 completes our reduction of the original updates (2) and (3) to the equivalent Algorithm 1. A runtime analysis comparing this efficient implementation (Algorithm 1) with a direct evaluation of the closed-form solutions in (4) and (5) is presented in Section 4.3.
At iteration 1 of Algorithm 1, both $q_{1}$ and $w^{(1)}_{\geq 2}$ depend on $\widetilde{X},X\in\mathbb{R}^{m\times N}$, yielding memory requirements of $\mathcal{O}(mN)$, often where $m\gg N$; for example, Llama3.1-8B requires over 30 GB just to store 128 samples of 2048-token sequences at float32, not including weight storage. Observing that OPTQ has memory requirements of $\mathcal{O}(N^{2})$, we can rewrite the first iteration of Algorithm 1 using only square matrices as below, where $G=\widetilde{X}^{T}X\in\mathbb{R}^{N\times N}$ and $H=\widetilde{X}^{T}\widetilde{X}\in\mathbb{R}^{N\times N}$; see Proposition A.2 for a justification.
$q_{1}$ $=\mathcal{Q}\left(\frac{G_{1,\geq 1}w-H_{1,\geq 2}w^{(0)}_{\geq 2}}{H_{11}}\right)$ (14)
$w^{(1)}_{\geq 2}$ $=(H_{\geq 2,\geq 2})^{-1}\left(G_{\geq 2,\geq 1}w-H_{\geq 2,1}q_{1}\right)$ (15)
Note that calculating $G$ and $H$ does not require storing $\widetilde{X},X$, as one can sequentially accumulate the outer products of each of the $m$ samples. Thus, this square matrix formulation reduces peak memory requirements to $\mathcal{O}(N^{2})$, yielding an 18$\times$ reduction in the case of Llama3.1-8B. We note that Colbert et al. (2024) similarly identify a memory optimization for GPFQ, but use singular value decompositions that may not scale well with $N$.
3.2 Theoretical Interpretation and Intuition
Theorem 3.1 and Lemma 3.2 connect the initial disciplined optimization formulation of Qronos, which alternates between (2) and (3), with its efficient practical implementation, described in Algorithm 1. These results guarantee that Qronos is simultaneously interpretable and scalable, explicitly correcting error from the mismatched input pairs $X$ and $\widetilde{X}$. Here, we provide deeper intuition in the context of previous state-of-the-art rounding algorithms, namely GPFQ and OPTQ.
GPFQ is a data-driven error-correcting PTQ method proposed in Lybrand and Saab (2021) and further analyzed in Zhang et al. (2023); Zhang and Saab (2023). When quantizing $w$, it interprets $Xw$ as the endpoint of the path $\sum_{j=1}^{t}w_{j}X_{j}$ for $t=1,...,N$, and handles mismatched inputs by aiming to match $\sum_{j=1}^{t}w_{j}X_{j}$ and $\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}$ for all $t$. More precisely, $q_{t}$ is selected as $\arg\min_{p\in\mathcal{A}}\|\sum_{j=1}^{t}w_{j}X_{j}-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-p\widetilde{X}_{t}\|_{2}^{2}$, as its name (greedy path following quantization) suggests.
Although path following handles the case when $X=\widetilde{X}$ well, additional considerations are required when $X\neq\widetilde{X}$ since, in such a case, the tails of the two paths generally do not align when $\sum_{i=t+1}^{N}w_{i}(X_{i}-\widetilde{X}_{i})\neq 0$. Qronos handles this drawback by adopting a natural remedy to replace the unquantized weights $w_{i},i\geq t+1$ by auxiliary weights $w^{(t)}_{i},i\geq t+1$ so that
$\sum_{i=1}^{t}q_{i}\widetilde{X}_{i}\ +\sum_{i=t+1}^{N}w^{(t)}_{i}\widetilde{X}_{i}\ \approx\ Xw=\sum_{i=1}^{N}w_{i}X_{i}.$
In fact, OPTQ (Frantar et al., 2023) explores a similar weight update idea, but only in the case where $X=\widetilde{X}$, by modifying the remaining unquantized weights after $q_{t}$ is selected. The Cholesky reformulation used in Algorithm 1 also resembles the key mechanism in OPTQ (Algorithm 2, omitting lazy batch updates and dampening, for simplicity). In this way, the runtime of Qronos scales similarly to OPTQ while also explicitly addressing the mismatch between $X$ and $\widetilde{X}$; see Section 4.3 for details. This unexpected connection of Qronos to OPTQ also allows us to derive a novel interpretation of OPTQ in Appendix D, which shows that it corrects the cumulative quantization error incurred over all the previous iterations. We believe this result is of independent interest, due to the widespread use of OPTQ. The proofs for all results are presented in the appendix.
4 Experiments
The core contribution of this work is Qronos—our principled data-driven rounding algorithm that alternates between (1) explicitly correcting quantization error due to both the weights and activations, and (2) diffusing excess error into future weights yet to be quantized. Thus, our primary comparison metric is preserving model quality in challenging quantization scenarios. We design our experiments to isolate the impact of the rounding function (stage 2 in Figure 1), while varying the quantization transforms (stage 1 in Figure 1).
Models & Datasets. We conduct experiments on the Llama3 (Grattafiori et al., 2024) model family using WikiText2 (Merity et al., 2016) for evaluation. We leverage the unmodified implementations and checkpoints made publicly available via the Huggingface libraries (Wolf et al., 2020). We use LightEval (Fourrier et al., 2023) to evaluate generalization via 5 zero-shot reasoning tasks: ARC (challenge and easy) (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), and Winogrande (Sakaguchi et al., 2021), and report the normalized average accuracy.
Setup. We implement Qronos in PyTorch (Paszke et al., 2019) using the Brevitas quantization library (Franco et al., 2025b), and quantize all models using a single AMD MI210 GPU with 64 GB of memory. Unless otherwise specified, we construct our calibration dataset using 128 random sequences of 2048 tokens sampled from the WikiText2 dataset for all data-driven PTQ algorithms. We compare Qronos against RTN and the unmodified Brevitas implementations of OPTQ and GPFQ, also leveraging the unmodified Brevitas implementations of the various quantization transforms.
Baselines. Our primary baselines are RTN, OPTQ, and GPFQ. For OPTQ, we use the standard dampened covariance matrix $\widetilde{H}=H+\lambda I$, where $\lambda$ is 1% of the average diagonal of $H$. We similarly use a dampened covariance matrix for Qronos, but instead choose $\lambda$ to be based on the maximum singular value of $H$ such that $\lambda=\alpha\cdot\sigma_{1}$, which limits the condition number of $\widetilde{H}$ to be less than $\alpha^{-1}$. We select $\alpha=1e^{-6}$ for weight-only quantization and $\alpha=1e^{-3}$ for weight-activation quantization. Additionally, we apply error correction block-by-block for OPTQ, GPFQ, and Qronos, such that $X=\widetilde{X}$ at the beginning of each block. Note that this corresponds precisely to the original OPTQ proposal (Frantar et al., 2023), and has become common practice for GPFQ (Colbert et al., 2024). Finally, when quantizing weights sequentially, we do so in descending order of the diagonals of $H$, as is also now common practice (IST-DASLab, 2022; Franco et al., 2025b).
Table 2: Weight-only quantization of Llama3 models. We individually apply various quantization transforms (stage 1) to isolate the impact of different rounding functions (stage 2) across bit widths.
    W4 W3
    WikiText2 ($\downarrow$) 0-shot ($\uparrow$) WikiText2 ($\downarrow$) 0-shot ($\uparrow$)
Stage 1 Stage 2     1B 3B 8B     1B 3B 8B     1B 3B 8B     1B 3B 8B
BF16 -     8.9 7.1 5.9     59.4 67.5 74.4     8.9 7.1 5.9     59.4 67.5 74.4
None RTN     18.0 10.1 8.4     49.1 60.8 67.4     2e4 1e4 3e4     32.3 32.4 32.6
OPTQ     10.4 7.8 6.5     54.3 63.4 70.2     42.5 13.8 11.4     37.5 48.1 53.8
GPFQ     10.4 7.8 6.5     56.0 65.2 71.2     35.3 13.4 11.1     35.7 49.9 53.5
Qronos     10.1 7.6 6.4     54.8 64.5 71.3     21.4 11.3 9.3     40.4 53.3 58.1
Smooth Quant RTN     15.2 9.6 8.1     51.4 61.5 67.5     6e3 9e3 5e4     32.7 32.9 31.4
OPTQ     10.4 7.9 6.6     56.2 65.3 70.2     29.6 13.6 12.6     37.0 46.9 47.3
GPFQ     10.8 7.9 6.7     53.9 64.4 69.9     30.1 14.7 12.9     36.5 44.8 45.4
Qronos     10.3 7.8 6.5     56.8 64.8 71.8     19.1 11.6 10.3     40.9 48.5 51.5
MagR RTN     13.8 10.3 7.2     53.3 57.9 69.5     2e3 2e3 5e4     33.7 33.4 35.2
OPTQ     10.3 8.0 6.5     56.2 60.3 69.7     19.8 13.2 8.3     45.5 49.7 59.2
GPFQ     10.4 8.5 6.5     55.8 58.9 70.6     25.0 14.4 8.3     42.3 47.7 60.6
Qronos     10.1 8.0 6.4     56.2 62.0 68.8     17.1 12.0 7.8     45.9 49.9 60.2
HIP RTN     13.8 8.8 7.2     52.0 62.8 70.0     7e2 3e2 1e2     34.2 33.3 36.3
OPTQ     9.9 7.6 6.3     56.8 66.1 72.1     16.1 10.3 8.8     44.1 56.6 59.2
GPFQ     9.9 7.6 6.3     56.5 65.7 72.0     16.6 10.4 8.6     44.9 54.8 60.5
Qronos     9.6 7.5 6.2     57.0 67.1 72.2     12.9 9.3 7.9     46.6 58.5 63.7
4.1 Weight-Only Quantization
We first present state-of-the-art 2-bit and 1.58-bit results for weight-only PTQ on Llama3, controlling for the quantization transform and grid selection while varying the rounding function. We quantize weights using the standard asymmetric weight quantizer (Frantar et al., 2023), where scaling factors and zero points are defined per-channel on the min-max grid. We combine Hadamard-based incoherence processing (HIP) (Tseng et al., 2024a; Ashkboos et al., 2024) with weight magnitude reduction (MagR) (Zhang et al., 2024) to jointly act as our quantization transform, as they are both known to be effective at 2-bit weight quantization (Chee et al., 2023; Adepu et al., 2024). We present our results in Table 1, as well as the BF16 baselines, and highlight that Qronos consistently outperforms existing rounding methods. For example, when compared to OPTQ, Qronos provides a 1.8$\times$ reduction in WikiText2 perplexity and $+$2.4% increase in average zero-shot accuracy for Llama3.2-1B at 2 bits, and an order of magnitude improvement in perplexity (45.3$\times$) at 1.58 bits.
Next, we present state-of-the-art 4-bit and 3-bit weight-only PTQ results (denoted W4 and W3, respectively) while independently demonstrating compatibility with 3 notable quantization transforms: SmoothQuant (Xiao et al., 2023), MagR, and HIP. Table 2 shows the results across three models in the Llama3 family. Qronos consistently provides higher quality quantized models than RTN, OPTQ, and GPFQ, as measured in both WikiText2 perplexity and average zero-shot accuracy. Consistent with emerging work on rotation-based quantization transforms (Chee et al., 2023; Tseng et al., 2024a), incoherence processing outperforms other transforms, with HIP + Qronos providing the best overall results. Note that HIP + OPTQ is similar in spirit to QuIP by Theorem 6 in (Chee et al., 2023), which equates LDLQ to OPTQ, with the notable difference that QuIP proposed random orthogonal matrices instead of Hadamard matrices.
4.2 Weight-Activation Quantization
We present 4-bit weight-activation quantization results with and without 4-bit KV cache quantization (denoted W4A4KV16 and W4A4KV4, respectively) while demonstrating compatibility with QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2025). Here, we quantize weights using the standard symmetric weight quantizer with per-channel scaling factors optimized via linear search over the mean square error loss between the full-precision and quantized weights. We quantize activations and KV caches using the standard dynamic asymmetric activation quantizer with per-token scaling factors and zero points defined on the min-max grid, as is common practice (Liu et al., 2025). When learning the rotations for SpinQuant, we do so via 100 iterations of Cayley SGD (Li et al., 2020) using 128 random samples from the C4 dataset (Raffel et al., 2020).
Table 3 shows the results across three models in the Llama3 family. Qronos again consistently outperforms RTN, OPTQ, and GPFQ as measured in both WikiText2 perplexity and average zero-shot accuracy. Consistent with emerging work on learned rotations (Liu et al., 2025; Hu et al., 2025; Franco et al., 2025a), SpinQuant outperforms QuaRot, with SpinQuant + Qronos providing the best overall results with and without KV cache quantization. We remark that our experiments use per-token quantization for both the activations and KV caches, while Ashkboos et al. (2024) and Liu et al. (2025) use per-group scaling for KV cache quantization.
Table 3: Weight-activation quantization of Llama3 models. We individually apply various quantization transforms (stage 1) to isolate the impact of different rounding functions (stage 2).
    W4A4KV16 W4A4KV4
    WikiText2 ($\downarrow$) 0-shot ($\uparrow$) WikiText2 ($\downarrow$) 0-shot ($\uparrow$)
Stage 1 Stage 2     1B 3B 8B     1B 3B 8B     1B 3B 8B     1B 3B 8B
BF16 -     8.9 7.1 5.9     59.4 67.5 74.4     8.9 7.1 5.9     59.4 67.5 74.4
QuaRot RTN     22.0 12.6 9.6     45.4 55.0 62.6     41.8 22.0 15.9     41.5 49.8 57.4
OPTQ     14.3 9.8 8.0     50.4 59.9 66.7     19.8 14.3 10.3     45.8 56.2 64.1
GPFQ     13.6 9.3 7.6     50.9 60.9 67.6     22.0 14.7 11.4     43.3 53.9 59.8
Qronos     12.9 9.2 7.4     51.0 61.7 68.0     18.0 11.6 9.3     47.1 57.3 65.1
SpinQuant RTN     13.8 9.7 7.9     50.4 61.9 67.8     17.5 12.2 9.8     49.3 58.5 64.8
OPTQ     13.5 9.3 7.7     52.2 61.1 68.3     17.4 11.8 9.3     48.8 58.6 64.9
GPFQ     13.4 9.2 7.6     51.1 61.7 66.8     21.1 14.3 10.9     44.7 53.3 60.9
Qronos     12.3 8.7 7.2     52.7 61.5 68.8     16.0 11.0 8.7     48.3 59.1 65.2
(a) Runtime of Rounding Algorithm
(b) Runtime of Quantization Pipeline
Figure 2: We compare the runtime of (a) the rounding algorithm and (b) the overall quantization pipeline as we scale the input features $N$, as measured on an AMD MI210. We average all measurements over 3 seeds and normalize to the runtime of OPTQ where $N=32$.
4.3 Runtime Analysis
Qronos acts as a drop-in replacement at stage 2 of the quantization pipeline shown in Figure 1. We remark that Qronos (and rounding algorithms in general) leave the compute graph unaltered, thus there is no inference overhead on top of that introduced by the quantization transform in stage 1. Since prior works already profile the downstream inference overhead of various compute graph alteration (e.g., Ashkboos et al. (2024) conclude that optimized Hadamard transforms add at most 7% overhead to Llama2 models), we focus our runtime analysis on the quantization pipeline.
We perform our runtime analysis using a single linear layer as our microbenchmark. We use a calibration set of $m=$10,000 random data sampled from a normal $K$-dimensional Gaussian distribution. The linear layer has $K\in[32,1024]$ inputs with $K/4$ outputs. Figure 2 shows how the runtime of OPTQ, GPFQ, and Qronos scale with $K$, where (a) isolates the algorithm runtime and (b) aggregates the quantization pipeline runtime. To highlight the benefits of our equivalent formulation, we implement a base version of Qronos that uses the iterates for $q_{t}$ and $w_{\geq t+1}^{(t)}$ from (4) and (5). Note that via Theorem 3.1 and Lemma 3.2, Algorithm 1 significantly improves the runtime scaling over the base version to match that of OPTQ, with a 13.8$\times$ reduction in algorithm runtime and a 3.6$\times$ reduction in overall runtime when $K=1024$. Compared with OPTQ, which only needs to collect $X$, GPFQ and Qronos require collecting both $\widetilde{X}$ and $X$ at each layer, which requires two forward passes (with and without quantization) and increases the overall quantization pipeline runtime. In practice, we observe that Qronos take 2$\times$ longer to quantize Llama3.2-1B when compared to OPTQ; however, from Meta (2024), quantization-aware training via QLoRA (Dettmers et al., 2023) took 1,300 GPU hours for Llama3.2-1B, while Qronos took 45 minutes on average.
5 Conclusions
We introduce Qronos—a new state-of-the-art rounding algorithm that alternates between correcting quantization error in both the weights and activations of previous layers and diffusing error into future weights within the current layer. Qronos is based on an interpretable and disciplined optimization framework, and it demonstrably subsumes and surpasses existing data-driven approaches, namely OPTQ and GPFQ. Our implementation exploits several optimizations that together yield orders of magnitude improvements in memory and compute efficiency. Our experiments isolate the impact of the rounding function in the quantization pipeline while varying transformations on a min-max grid. Our results demonstrate that Qronos consistently offers improvements over previous state-of-the-art methods when quantizing weights, activations, and/or KV caches to 4 bits or fewer. That said, our results are intentionally limited to the min-max quantization grid to focus our experiments on transformations and rounding; we believe our results could be further improved by leveraging weight and activation distributions to design quantization grids that are more effective than the min-max grid used in this work, possibly with non-uniform grids via vector quantization.
6 Ethics Statement
Our work focuses on efficient post-training quantization (PTQ) algorithms aimed at reducing the computational and memory demands of neural networks, both in terms of GPU usage and storage, and without requiring any retraining. Our algorithm is highly efficient and applicable to large models. By applying our method, energy efficiency can be improved, and deployment on resource-constrained edge devices may become feasible. This may help broaden access to powerful models and reduce the environmental impact of model deployment.
However, as with other compression techniques, our method could also be used to deploy powerful models in cases where fairness, privacy, or safety are not sufficiently considered. We encourage the community to evaluate compressed models for unintended consequences in downstream applications. Our work uses only publicly available datasets and models, and does not involve human subjects, private data, or personally identifiable information.
7 Reproducibility Statement
We intend to open source our code and instructions on how to reproduce the various experiments, including random seeds. The code will also point to all datasets and models used for evaluation.
Acknowledgment
We gratefully acknowledge partial support by National Science Foundation, via the DMS-2410717 grant.
Appendix A Preliminary Propositions
Proposition A.1.
The update rule given by
$q_{t}$ $=\underset{p\in\mathcal{A}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-p\widetilde{X}_{t}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j}\|^{2},$
$w^{(t)}_{\geq t+1}$ $=\underset{(v_{t+1},...,v_{N})\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}v_{j}\widetilde{X}_{j}\|^{2}.$
has closed-form expressions
$q_{t}=\mathcal{Q}\left(\frac{\langle Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j},\widetilde{X}_{t}\rangle}{\|\widetilde{X}_{t}\|^{2}}\right)$
and
$w^{(t)}_{\geq t+1}=\widetilde{X}_{\geq t+1}^{\dagger}\left(Xw-\widetilde{X}_{\leq t}q_{\leq t}\right).$
Proof.
For $q_{t}$, the corresponding optimization objective function is a one-dimensional quadratic function of $p$. Since minimizing a quadratic function over a discrete set $\mathcal{A}$ reduces to rounding its real-valued minimizer, we compute the real-valued minimizer
$\frac{\langle Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j},\widetilde{X}_{t}\rangle}{\|\widetilde{X}_{t}\|^{2}}.$
Thus, we obtain the closed-form expression of $q_{t}$,
$q_{t}=\mathcal{Q}\left(\frac{\langle Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j},\widetilde{X}_{t}\rangle}{\|\widetilde{X}_{t}\|^{2}}\right),$
where $\mathcal{Q}$ is the round-to-nearest operator.
For $w^{(t)}_{\geq t+1}$, the corresponding optimization problem is an unconstrained least-square problem in the form of $\min_{v\in\mathbb{R}^{N-t}}\|Ax-b\|^{2}$, with $A=\widetilde{X}_{\geq t+1}$ and $b=Xw-\widetilde{X}_{\leq t}q_{\leq t}$. Thus, the minimizer is given by $A^{\dagger}b$, which gives the desired closed-form expression. ∎
Proposition A.2.
The update rule given by
$q_{1}=\mathcal{Q}\left(\dfrac{\widetilde{X}_{1}^{\top}(Xw-\widetilde{X}_{\geq 2}w^{(0)}_{\geq 2})}{\|\widetilde{X}_{1}\|^{2}}\right),$
$w^{(1)}_{\geq 2}=\widetilde{X}_{\geq 2}^{\dagger}\left(Xw-\widetilde{X}_{1}q_{1}\right)$
is equivalent to
$q_{1}$ $=\mathcal{Q}\left(\frac{G_{1,\geq 1}w-H_{1,\geq 2}w^{(0)}_{\geq 2}}{H_{11}}\right)$
$w^{(1)}_{\geq 2}$ $=(H_{\geq 2,\geq 2})^{-1}\left(G_{\geq 2,\geq 1}w-H_{\geq 2,1}q_{1}\right),$
where $G=\widetilde{X}^{T}X\in\mathbb{R}^{N\times N}$ and $H=\widetilde{X}^{T}\widetilde{X}\in\mathbb{R}^{N\times N}$.
Proof.
For $q_{1}$, we have $\widetilde{X}_{1}^{\top}X=(\widetilde{X}^{\top}X)_{1,\geq 1}=G_{1,\geq 1}$. Also, $\widetilde{X}_{1}^{\top}\widetilde{X}_{\geq 2}=(\widetilde{X}^{\top}\widetilde{X})_{1,\geq 2}=H_{1,\geq 2}$. Thus, $\widetilde{X}_{1}^{\top}(Xw-\widetilde{X}_{\geq 2}w^{(0)}_{\geq 2})=G_{1,\geq 1}w-H_{1,\geq 2}w^{(0)}_{\geq 2}$. Further, $\|\widetilde{X}_{1}\|^{2}=(\widetilde{X}^{\top}\widetilde{X})_{11}=H_{11}$. This gives the equivalence for updating $q_{1}$.
For $w^{(1)}_{\geq 2}$, $\widetilde{X}_{\geq 2}$ is given by $(\widetilde{X}_{\geq 2}^{\top}\widetilde{X}_{\geq 2})^{-1}\widetilde{X}_{\geq 2}^{\top}=(H_{\geq 2,\geq 2})^{-1}\widetilde{X}_{\geq 2}^{\top}$. Then
$\widetilde{X}_{\geq 2}^{\dagger}\left(Xw-\widetilde{X}_{1}q_{1}\right)$ $=(H_{\geq 2,\geq 2})^{-1}\widetilde{X}_{\geq 2}^{\top}\left(Xw-\widetilde{X}_{1}q_{1}\right)$
$=(H_{\geq 2,\geq 2})^{-1}\left((\widetilde{X}^{\top}X)_{\geq 2,\geq 1}w-(\widetilde{X}^{\top}\widetilde{X})_{\geq 2,1}q_{1}\right)$
$=(H_{\geq 2,\geq 2})^{-1}\left(G_{\geq 2,\geq 1}w-H_{\geq 2,1}q_{1}\right).$
This gives the equivalence for updating $w^{(1)}_{\geq 2}.$
Appendix B Proof of Theorem 3.1
Proof.
We use induction to prove the theorem. Since at $t=1$ equations (2), (3) and equations (6), (7) are identical, the base case is trivially true. Now we proceed with the induction, assuming $\hat{w}^{(t)}_{\geq t+1}=w^{(t)}_{\geq t+1}$ and $\hat{q}_{t}=q_{t}$.
Using definition (2) and Proposition A.1, we can obtain the closed-form expression,
$q_{t+1}=\mathcal{Q}\left(\frac{\langle Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+2}^{N}w^{(t)}_{j}\widetilde{X}_{j},\widetilde{X}_{t+1}\rangle}{\|\widetilde{X}_{t+1}\|^{2}}\right),$
where $\mathcal{Q}$ is the RTN operator. Next we note that (3), which is used to compute $w^{(t)}_{\geq t+1}$, implies that $Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j}$ is orthogonal to the column space of $\widetilde{X}_{\geq t+1}$. This in turn implies that $\langle Xw-\sum_{j=1}^{t}q_{j}\tilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j},\widetilde{X}_{t+1}\rangle=0$. Then we can compute,
$q_{t+1}$ $=\mathcal{Q}\left(\frac{\langle Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+2}^{N}w^{(t)}_{j}\widetilde{X}_{j},\widetilde{X}_{t+1}\rangle}{\|\widetilde{X}_{t+1}\|^{2}}\right)$
$=\mathcal{Q}\left(\frac{\langle Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j}+w^{(t)}_{t+1}\widetilde{X}_{t+1},\widetilde{X}_{t+1}\rangle}{\|\widetilde{X}_{t+1}\|^{2}}\right)$
$=\mathcal{Q}\left(\frac{\langle w^{(t)}_{t+1}\widetilde{X}_{t+1},\widetilde{X}_{t+1}\rangle}{\|\widetilde{X}_{t+1}\|^{2}}\right)$
$=\mathcal{Q}\left(w^{(t)}_{t+1}\right)=\mathcal{Q}\left(\hat{w}^{(t)}_{t+1}\right)=\hat{q}_{t+1},$
where in the last two inequalities, we used the induction hypothesis $\hat{w}^{(t)}_{\geq t+1}=w^{(t)}_{\geq t+1}$ and the update rule (8).
Next, we prove $\hat{w}^{(t+1)}_{\geq t+2}=w^{(t+1)}_{\geq t+2}.$ We first compute
$w^{(t+1)}_{\geq t+2}$ $=\underset{v_{\geq t+2}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t+1}q_{j}\widetilde{X}_{j}-\sum_{j=t+2}^{N}v_{j}\widetilde{X}_{j}\|^{2}$
$=\underset{v_{\geq t+2}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j}+(w^{(t)}_{t+1}-q_{t+1})\widetilde{X}_{t+1}+\sum_{j=t+2}^{N}(w^{(t)}_{j}-v_{j})\widetilde{X}_{j}\|^{2}.$
Due to the update rule (3), $Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j}$ is orthogonal to the column span of $\widetilde{X}_{\geq t+1}$, hence to $(w^{(t)}_{t+1}-q_{t+1})\widetilde{X}_{t+1}+\sum_{j=t+2}^{N}(w^{(t)}_{j}-v_{j})\widetilde{X}_{j}$. Then, we have
$w^{(t+1)}_{\geq t+2}$ $=\underset{v_{\geq t+2}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j}+(w^{(t)}_{t+1}-q_{t+1})\widetilde{X}_{t+1}+\sum_{j=t+2}^{N}(w^{(t)}_{j}-v_{j})\widetilde{X}_{j}\|^{2}$
$=\underset{v_{\geq t+2}}{\mathrm{argmin}}\frac{1}{2}\|(\hat{w}^{(t)}_{t+1}-\hat{q}_{t+1})\widetilde{X}_{t+1}+\sum_{j=t+2}^{N}(\hat{w}^{(t)}_{j}-v_{j})\widetilde{X}_{j}\|^{2}=\hat{w}^{(t+1)}_{\geq t+2},$
where we used the Pythagorean theorem, the induction hypothesis $\hat{w}^{(t)}_{\geq t+1}=w^{(t)}_{\geq t+1}$, and the fact $q_{t+1}=\hat{q}_{t+1}$. This completes the induction. ∎
Appendix C Proof of Lemma 3.2
Throughout this section, we denote $H_{\geq t,\geq t}=X_{\geq t}^{\top}X_{\geq t}\in\mathbb{R}^{(N-t+1)\times(N-t+1)}$ and $H^{-1}_{\geq t,\geq t}=(X_{\geq t}^{\top}X_{\geq t})^{-1}\in\mathbb{R}^{(N-t+1)\times(N-t+1)}$. We will begin with a few preliminary lemmas before we prove Lemma 3.2. While some of these lemmas may already be known, we are not aware of any rigorous proofs in the literature. Thus, we provide our proofs here for completeness.
Lemma C.1.
Denote by $[H^{-1}_{\geq t,\geq t}]_{11}$ the first entry of $H^{-1}_{\geq t,\geq t}$ and by $[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}\in\mathbb{R}^{N-t}$ the first column of $H^{-1}_{\geq t,\geq t}$ albeit with the first entry removed. Then
$(X_{\geq t+1}^{\top}X_{\geq t+1})^{-1}X_{\geq t+1}^{\top}X_{t}=-\frac{[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}}{[H^{-1}_{\geq t,\geq t}]_{11}}.$
Proof.
[…]
$(X_{\geq t+1}^{\top}X_{\geq t+1})^{-1}X_{\geq t+1}^{\top}X_{t}=-\frac{\mathbf{b}}{r}$
and finishes the proof. ∎
The next lemma establishes how one can efficiently compute $H^{-1}_{\geq t+1,\geq t+1}$ from $H^{-1}_{\geq t,\geq t}$.
Lemma C.2.
$H^{-1}_{\geq t+1,\geq t+1}$ can be efficiently computed from $H^{-1}_{\geq t,\geq t}$ via
$H^{-1}_{\geq t+1,\geq t+1}=\left(H^{-1}_{\geq t,\geq t}-\frac{1}{[H^{-1}_{\geq t,\geq t}]_{11}}[H^{-1}_{\geq t,\geq t}]_{\geq 1,1}[H^{-1}_{\geq t,\geq t}]_{1,\geq 1}\right)_{\geq 2,\geq 2}.$
We note that this is a simple rank-$1$ update followed by a submatrix slicing.
Proof.
We first recall a more general inverse formula for $2\times 2$ block matrix using the Schur complement. Consider the $2\times 2$ block matrix
$M=\begin{pmatrix}A&B\\ C&D\end{pmatrix}.$
When $A$ is invertible, the inverse of $M$ is given by
$M^{-1}=\begin{pmatrix}A^{-1}+A^{-1}BS^{-1}CA^{-1}&-A^{-1}BS^{-1}\\ -S^{-1}CA^{-1}&S^{-1}\end{pmatrix},$ (16)
where $S=D-CA^{-1}B$ is the Schur complement of $A$ in $M$.
When $A$ is a scalar $a$ and ${M}$ is symmetric, i.e.
${M}=\begin{pmatrix}a&b^{\top}\\ b&D\end{pmatrix},$
this formula becomes
${M}^{-1}=\begin{pmatrix}a^{-1}+a^{-2}b^{\top}S^{-1}b&-a^{-1}b^{\top}S^{-1}\\ -a^{-1}S^{-1}b&S^{-1}\end{pmatrix},$
where $S=D-a^{-1}bb^{\top}$.
By the Sherman–Morrison formula [Horn and Johnson, 2012], we have
$D^{-1}$ $=S^{-1}-\frac{S^{-1}bb^{\top}S^{-1}}{a+b^{\top}S^{-1}b}$
$=S^{-1}-\frac{a^{-2}S^{-1}bb^{\top}S^{-1}}{a^{-1}+a^{-2}b^{\top}S^{-1}b}.$
Returning to our setting where ${M}^{-1}=H^{-1}_{\geq t,\geq t}$ and ${D}^{-1}=H^{-1}_{\geq t+1,\geq t+1}$, we have
$H^{-1}_{\geq t+1,\geq t+1}$ $=[H^{-1}_{\geq t,\geq t}]_{\geq 2,\geq 2}-\frac{1}{[H^{-1}_{\geq t,\geq t}]_{11}}[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}[H^{-1}_{\geq t,\geq t}]_{1,\geq 2}$
$=[H^{-1}_{\geq t,\geq t}]_{\geq 2,\geq 2}-\frac{1}{[H^{-1}_{\geq t,\geq t}]_{11}}\left([H^{-1}_{\geq t,\geq t}]_{\geq 1,1}[H^{-1}_{\geq t,\geq t}]_{1,\geq 1}\right)_{\geq 2,\geq 2}$
$=\left(H^{-1}_{\geq t,\geq t}-\frac{1}{[H^{-1}_{\geq t,\geq t}]_{11}}[H^{-1}_{\geq t,\geq t}]_{\geq 1,1}[H^{-1}_{\geq t,\geq t}]_{1,\geq 1}\right)_{\geq 2,\geq 2}.$
Using the above lemma and Cholesky decomposition [Horn and Johnson, 2012], we can further simplify the right hand side in Lemma C.1 via the following lemma.
Lemma C.3.
Let $H^{-1}=(X^{\top}X)^{-1}$ and $H^{-1}=LL^{\top}$ be its Cholesky decomposition where $L$ is a lower triangular matrix, then
$\frac{[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}}{[H^{-1}_{\geq t,\geq t}]_{11}}=\frac{L_{\geq t+1,t}}{L_{tt}}\in\mathbb{R}^{N-t}$
holds for all $t\in[N-1]$.
Proof.
We first prove that given the Cholesky decomposition $H^{-1}=LL^{\top}$, the Cholesky decomposition of $H^{-1}_{\geq t,\geq t}$ is $H^{-1}_{\geq t,\geq t}=(L_{\geq t,\geq t})(L_{\geq t,\geq t})^{\top}$ for all $t\in[N]$, where $H^{-1}_{\geq t,\geq t}=(X_{\geq t}^{\top}X_{\geq t})^{-1}\in\mathbb{R}^{(N-t+1)\times(N-t+1)}$.
Let us proceed by induction. The base-case when $t=1$ holds by assumption, and we now assume the result holds for $t$. By Lemma C.2, the updated inverse Hessian $H^{-1}_{\geq t+1,\geq t+1}=\left(H^{-1}_{\geq t,\geq t}-\frac{1}{[H^{-1}_{\geq t,\geq t}]_{11}}[H^{-1}_{\geq t,\geq t}]_{\geq 1,1}[H^{-1}_{\geq t,\geq t}]_{1,\geq 1}\right)_{\geq 2,\geq 2}$. Thus,
$\Big{(}(L_{\geq t,\geq t})$ $(L_{\geq t,\geq t})^{\top}-\frac{1}{L_{tt}^{2}}((L_{\geq t,\geq t})_{11}\cdot[L_{\geq t,\geq t}]_{\geq 1,1})((L_{\geq t,\geq t})_{11}\cdot[L_{\geq t,\geq t}]_{\geq 1,1})^{\top}\Big{)}_{\geq 2,\geq 2}$
$=$ $\left((L_{\geq t,\geq t})(L_{\geq t,\geq t})^{\top}-[L_{\geq t,\geq t}]_{\geq 1,1}[L_{\geq t,\geq t}]_{\geq 1,1}^{\top}\right)_{\geq 2,\geq 2}$
$=$ $((L_{\geq t,\geq t})_{\geq 2,\geq 2})((L_{\geq t,\geq t})_{\geq 2,\geq 2})^{\top}$
$=$ $(L_{\geq t+1,\geq t+1})(L_{\geq t+1,\geq t+1})^{\top}$
This finishes the induction and we have Cholesky decomposition $H^{-1}_{\geq t,\geq t}=(L_{\geq t,\geq t})(L_{\geq t,\geq t})^{\top}$ for all $t\in[N]$. To finish the proof, let $M=RR^{\top}$ be the Cholesky decomposition of any positive definite matrix $M$. By a direct computation, the first column of $M$ is $R[R^{\top}]_{\geq 1,1}=R_{11}\cdot[R]_{\geq 1,1}$ and the first entry $M_{11}=R_{11}^{2}$. Then we have $\frac{M_{\geq 1,1}}{M_{11}}=\frac{[R]_{\geq 1,1}}{R_{11}}$ which implies that $\frac{M_{\geq 2,1}}{M_{11}}=\frac{[R]_{\geq 2,1}}{R_{11}}$. In our case, we have $H^{-1}_{\geq t,\geq t}=(L_{\geq t,\geq t})(L_{\geq t,\geq t})^{\top}$ in the place of $M=RR^{\top}$. Thus,
$\frac{[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}}{[H^{-1}_{\geq t,\geq t}]_{11}}=\frac{[L_{\geq t,\geq t}]_{\geq 2,1}}{[L_{\geq t,\geq t}]_{11}}=\frac{L_{\geq t+1,t}}{L_{tt}}.$
With the above preliminary lemmas, now we are ready to prove Lemma 3.2
Proof of Lemma 3.2.
Since we initialize with $w^{(0)}=w$, $q_{1}=\mathcal{Q}(w_{1})$ always holds. Thus the two iterations produce the same $q_{1}$ and $w^{(0)}_{\geq 1}$ We proceed by induction. Assume at step $t$ that $q_{t}$ and $w^{(t-1)}_{\geq t}$ resulting from the update rules (10) and (11) match those following update rules (12) and (13). In order to complete the induction, it suffices to show that (11) and (13) produce the same $w^{(t)}_{\geq t+1}$, which naturally results in the same $q_{t+1}=\mathcal{Q}(w^{(t)}_{t+1})$.
To that end, we note that the optimization problem defined by (11) has a unique least-square solution as $X_{\geq t+1}$ has full column rank. The minimizer is given by
${w}_{\geq t+1}^{(t)}$ $=w_{\geq t+1}^{(t-1)}+(w^{(t-1)}_{t}-q_{t})X_{t+1:}^{\dagger}X_{t}$
$=w_{\geq t+1}+(w^{(t-1)}_{t}-q_{t})(X_{\geq t+1}^{\top}X_{\geq t+1})^{-1}X_{\geq t+1}^{\top}X_{t}$
By Lemma C.1, we have
$(X_{\geq t+1}^{\top}X_{\geq t+1})^{-1}X_{\geq t+1}^{\top}X_{t}=-\frac{[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}}{[H^{-1}_{\geq t,\geq t}]_{11}}.$
Lastly, Lemma C.3 gives us
$\frac{[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}}{[H^{-1}_{\geq t,\geq t}]_{11}}=\frac{L_{\geq t+1,t}}{L_{tt}}\in\mathbb{R}^{N-t}.$
This matches $\Delta_{t+1}$ in (13) and completes our induction. ∎
Appendix D A Novel Interpretation of OPTQ
Algorithm 2 OPTQ: Quantize a layer $W$ given inverse Hessian $H^{-1}=(X^{\top}X)^{-1}$.
1:for every $w$ in $W$ in parallel do
2:   $q=\mathbf{0}^{N}$$\triangleright$ Initialize quantized neuron
3:   $H^{-1}=LL^{\top}$triangleright Perform Cholesky decomposition
4:   for $t=1$ to $N$ do $\triangleright$ Iterate over rows
5:     $q_{t}=\mathcal{Q}(w_{t})$
6:     $w_{\geq t}\leftarrow w_{\geq t}-L_{\geq t,t}\cdot(w_{t}-q_{t})/L_{tt}$$\triangleright$ Update remaining weights
7:   end for
8:end for
9:return $Q$
For our final result of this paper, we observe that updates of $w^{(t)}_{\geq t+1}$ via (11) can be interpreted by observing that the term $(q_{t}-w^{(t-1)}_{t})X_{t}$ represents the error introduced by quantizing $w^{(t-1)}_{t}$. The optimization problem (11) seeks to mitigate this error by adjusting future weights so as to minimize the resulting distortion, measured in the $\ell_{2}$-norm. Notably, this step does not explicitly attempt to correct errors introduced by earlier quantization steps $11,\dots,t-$ 1. However, by combining the proof of Theorem 3.1 in the case when $X=\widetilde{X}$ with Lemma 3.2, we arrive at Corollary D.1, which provides a novel interpretation of OPTQ. It shows—perhaps unexpectedly—that Algorithm 2 optimally corrects the cumulative quantization error incurred over the first $t$ entries of $w$.
Corollary D.1.
The OPTQ update rules from Algorithm 2 are equivalent to
$q_{t}$ $=\underset{p\in\mathcal{A}}{\arg\min}\,\frac{1}{2}\|Xw-\sum_{j=1}^{t-1}q_{j}X_{j}-pX_{t}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}X_{j}\|^{2},$ (17)
$w^{(t)}_{\geq t+1}$ $=\underset{(v_{t+1},\dots,v_{N})\in\mathbb{R}^{N-t}}{\arg\min}\,\frac{1}{2}\|Xw-\sum_{j=1}^{t}q_{j}X_{j}-\sum_{j=t+1}^{N}v_{j}X_{j}\|^{2},$ (18)
with $w^{(0)}_{\geq 1}=w$. In other words, the updated weights and quantized weights at every iteration $t$ that are produced by Algorithm 2 are identical to those produced by equations (17) and (18). In particular, (18) shows that, at each step, the updated OPTQ weights $w^{(t)}_{\geq t+1}$indeed optimally correct for the errors produced by the hitherto quantized sequence $q_{1},...,q_{t}$.
Proof.
The proof is based on induction on both arguments of the trajectory. Let $\{(\hat{w}^{(t-1)}_{\geq t},\hat{q}_{t})\}_{t=1}^{N}$ denote the trajectory generated by update rules (17), (18). And let $\{(w^{(t-1)}_{\geq t},q_{t})\}_{t=1}^{N}$ be the trajectory generated by Algorithm 2. Our goal is to prove $(\hat{w}^{(t-1)}_{\geq t},\hat{q}_{t})=(w^{(t-1)}_{\geq t},q_{t})$ for $t=1,\dots,N$.
By Lemma 3.2, the trajectory $\{(w^{(t-1)}_{\geq t},q_{t})\}_{t=1}^{N}$ generated using Cholesky decomposition in Algorithm 2 can be equivalently regarded as generated from (10), (11). Thus, we will use (10), (11) as the update rule of $w^{(t-1)}_{\geq t}$ and $q_{t}$ in the rest of our proof. In the base case, $\hat{w}^{(0)}_{\geq 1}=w^{(0)}_{\geq 1}$ are both initialized with $w$ and
$\hat{q}_{1}=\underset{p\in\mathcal{A}}{\mathrm{argmin}}\frac{1}{2}\|Xw-pX_{1}-\sum_{j=2}^{N}w^{(0)}_{j}X_{j}\|^{2}=\underset{p\in\mathcal{A}}{\mathrm{argmin}}\frac{1}{2}\|(w_{1}-p)X_{1}\|^{2}=\mathcal{Q}(w_{1})=q_{1}.$
Thus $(w^{(0)}_{\geq 1},q_{1})=(\hat{w}^{(0)}_{\geq 1},\hat{q}_{1})$. Assume $(\hat{w}^{(t-1)}_{\geq t},\hat{q}_{t})=(w^{(t-1)}_{\geq t},q_{t})$ holds true. Now we proceed to prove $(\hat{w}^{(t)}_{\geq t+1},\hat{q}_{t+1})=(w^{(t)}_{\geq t+1},q_{t+1})$.
Step 1: We first prove $\hat{w}^{(t)}_{\geq t+1}={w}^{(t)}_{\geq t+1}$. By construction,
$\hat{w}^{(t)}_{\geq t+1}=\underset{v_{\geq t+1}\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}v_{j}X_{j}\|^{2}.$
For an arbitrary $v_{\geq t+1}\in\mathbb{R}^{N-t}$,
$Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}$ $-\sum_{j=t+1}^{N}v_{j}X_{j}=$
$\underbrace{(Xw-\sum_{j=1}^{t-1}\hat{q}_{j}X_{j}-\sum_{j=t}^{N}\hat{w}^{(t-1)}_{j}X_{j})}_{\textbf{(I)}}+\underbrace{\left((\hat{w}^{(t-1)}_{t}-\hat{q}_{t})X_{t}+\sum_{j=t+1}^{N}(\hat{w}^{(t-1)}_{j}-v_{j})X_{j}\right)}_{\textbf{(II)}}.$
[…]
$\left\|Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}v_{j}X_{j}\right\|^{2}=\left\|\textbf{(I)}\right\|^{2}+\left\|\textbf{(II)}\right\|^{2}.$
Notice that (I) does not depend on $v_{\geq t+1}$. Furthermore, $\hat{w}^{(t-1)}_{\geq t+1}$ and $\hat{q}_{t}$ in (II) can be replaced by ${w}^{(t-1)}_{\geq t+1}$ and $q_{t}$ respectively using our induction hypothesis. Thus,
$\hat{w}^{(t)}_{\geq t+1}$ $=\underset{v_{\geq t+1}\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}v_{j}X_{j}\|^{2}$
$=\underset{v_{\geq t+1}\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|(\hat{w}^{(t-1)}_{t}-\hat{q}_{t})X_{t}+\sum_{j=t+1}^{N}(\hat{w}^{(t-1)}_{j}-v_{j})X_{j}\|^{2}$
$=\underset{v_{\geq t+1}\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|({w}^{(t-1)}_{t}-{q}_{t})X_{t}+\sum_{j=t+1}^{N}({w}^{(t-1)}_{j}-v_{j})X_{j}\|^{2}$
$={w}^{(t)}_{\geq t+1}.$
Step 2: Now we prove $\hat{q}_{t+1}=q_{t+1}$. We just constructed
$\hat{w}^{(t)}_{\geq t+1}=\underset{v_{\geq t+1}\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}v_{j}X_{j}\|^{2}.$
This implies
$Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}\hat{w}^{(t)}_{j}X_{j}=P_{X_{\geq t+1}^{\perp}}(Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j})\in X_{\geq t+1}^{\perp}.$ (19)
By construction, we have
$\hat{q}_{t+1}$ $=\underset{q\in\mathcal{A}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-qX_{t+1}-\sum_{j=t+2}^{N}\hat{w}^{(t)}_{j}X_{j}\|^{2}$
$=\mathcal{Q}\left(\frac{\langle X_{t+1},Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+2}^{N}\hat{w}^{(t)}_{j}X_{j}\rangle}{\|X_{t+1}\|^{2}}\right).$
Then we can use (19) to deduce
$\frac{\langle X_{t+1},Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+2}^{N}\hat{w}^{(t)}_{j}X_{j}\rangle}{\|X_{t+1}\|^{2}}$
$=\frac{\langle X_{t+1},Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}\hat{w}^{(t)}_{j}X_{j}+X_{t+1}\hat{w}^{(t)}_{t+1}\rangle}{\|X_{t+1}\|^{2}}$
$=\frac{\langle X_{t+1},Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}\hat{w}^{(t)}_{j}X_{j}\rangle}{\|X_{t+1}\|^{2}}+\frac{\langle X_{t+1},X_{t+1}\hat{w}^{(t)}_{t+1}\rangle}{\|X_{t+1}\|^{2}}$
$=\frac{\langle X_{t+1},X_{t+1}\hat{w}^{(t)}_{t+1}\rangle}{\|X_{t+1}\|^{2}}$
$=\hat{w}^{(t)}_{t+1}$
$=w^{(t)}_{t+1}.$
The last step $\hat{w}^{(t)}_{t+1}=w^{(t)}_{t+1}$ follows from what we just proved in Step 1 that $\hat{w}^{(t)}_{\geq t+1}={w}^{(t)}_{\geq t+1}$. Thus we know
$\hat{q}_{t+1}=\mathcal{Q}(\hat{w}^{(t)}_{t+1})=\mathcal{Q}(w^{(t)}_{t+1})=q_{t+1}.$
This completes our induction. ∎