### v2

# Qronos: Correcting the Past by Shaping the Future… in Post-Training Quantization

Abstract
We introduce Qronos—a new state-of-the-art post-training quantization algorithm that sequentially rounds and updates neural network weights. Qronos not only explicitly corrects errors due to both weight and activation quantization, but also errors resulting from quantizing previous layers. Our iterative algorithm is based on an interpretable and disciplined optimization framework that subsumes and surpasses existing data-driven approaches. At each step, Qronos alternates between error correction and diffusion via optimal update rules. Importantly, we prove that Qronos admits an efficient implementation that uses the Cholesky decomposition for solving least-squares problems. We integrate our implementation into the open-source Brevitas quantization library [^1] https://github.com/i-colbert/brevitas/tree/qronos/src/brevitas_examples/llm and demonstrate compatibility with existing transformation techniques such as Hadamard-based incoherence processing and weight-activation scaling equalization, among others. We evaluate Qronos using recent autoregressive language generation models in the Llama3 family; Qronos consistently outperforms previous state-of-the-art adaptive rounding methods when quantizing the weights, activations, and/or KV caches to 4 bits or fewer.
1 Introduction
Recent advances in post-training quantization (PTQ) have enabled the practical use of few-bit weights and activations for large language model (LLM) inference, typically by focusing on one or both aspects of the quantization pipeline, visualized in Figure 1. The first aspect involves modifying the weights and activations of a model to make them more amenable to quantization, often through transformations that exploit invariances within the compute graph. The second aspect more directly concerns the design of the quantization mapping itself; it involves using data to minimize quantization error by either calibrating the quantization grid—defined by a bit width, scaling factor, and zero point—or adaptively rounding the (potentially transformed) weights.
The latest innovations in PTQ, including (Ashkboos et al., 2024; Liu et al., 2025), among many others, are skewed towards proposing and improving transformations that address the quantization challenges exacerbated in LLMs. These studies often only consider round-to-nearest (RTN) and OPTQ (Frantar et al., 2023), also known as GPTQ. Meanwhile, our work explicitly focuses on improving the rounding method while remaining compatible with these transformations to ultimately yield quantized models that more closely resemble their high-precision counterparts. We are in part motivated by the latest data-driven rounding algorithms, namely OPTQ (Frantar et al., 2023) and GPFQ (Lybrand and Saab, 2021). OPTQ, which is widely used in the literature, is a sequential algorithm that alternates between rounding the current weight and adjusting future unquantized weights to reduce error, a process we refer to as diffusion. However, OPTQ does not correct for quantization error in the activations or for error introduced by quantizing previous layers. In contrast, GPFQ explicitly corrects for quantization error in both the weights and activations in each rounding iteration, as well as error from previous layers, but does not adjust the future unquantized weights.
Contributions. We introduce Qronos—a new state-of-the-art post-training quantization algorithm that sequentially rounds and updates neural network weights. Qronos demonstrably subsumes and surpasses OPTQ and GPFQ, explicitly correcting quantization error in both the weights and activations of previous layers while diffusing error into future weights. In Section 3, we start by presenting Qronos in a well-disciplined and mathematically interpretable form, then rigorously derive an equivalent computationally efficient implementation. Our implementation exploits existing optimizations for OPTQ and GPFQ, such as Cholesky decomposition and block-level error diffusion. Indeed, in Section 4.3, we show our reductions significantly improve algorithm scaling, yielding up to a 13.8$\times$ improvement in algorithm runtime for a single-layer microbenchmark. As a theoretical by-product, we also derive a novel interpretation of OPTQ in Appendix F, which shows that it greedily corrects the cumulative quantization error incurred over previous iterations.
We empirically evaluate Qronos using the Llama3 model family (Grattafiori et al., 2024) and compare against RTN, OPTQ, and GPFQ while demonstrating compatibility with notable transformations for both weight-only quantization (e.g., SmoothQuant (Xiao et al., 2023) and incoherence processing Chee et al. (2023)) and weight-activation quantization (e.g., QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2025)). To our knowledge, our work is the first to isolate the impact of the rounding algorithm using a fixed quantization grid for each transformation method (or lack thereof).
Our experiments show that Qronos consistently yields marked improvement over existing rounding methods; for example, when quantizing the weights of Llama3 models to 2 bits or fewer, as shown in Table 1, Qronos provides massive improvements in perplexity over RTN, OPTQ, and GPFQ.
Table 1: Weight-only quantization of Llama3 models. We jointly apply Hadamard-based incoherence processing (Ashkboos et al., 2024) and weight magnitude reduction (Zhang et al., 2024) as quantization transforms (stage 1 in Figure 1) and compare different rounding functions (stage 2).
WikiText2 ($\downarrow$) 0-shot ($\uparrow$)
1B 3B 8B     1B 3B 8B
BF16 - 8.9 7.1 5.9     59.4 67.5 74.4
2-bit RTN 3e3 5e3 3e3     32.4 32.2 33.0
OPTQ 24.6 13.2 10.4     39.3 47.3 55.2
GPFQ 25.8 14.4 11.3     38.6 46.9 51.8
Qronos 17.8 11.4 9.3     42.6 50.7 55.8
1.58-bit RTN 5e5 4e4 9e4     32.3 32.9 32.2
OPTQ 2e2 52.0 43.3     32.7 32.5 34.9
GPFQ 1e2 51.3 35.8     32.4 32.6 33.4
Qronos 39.3 22.8 18.0     34.8 36.5 37.8
2 Background and Related Work
Figure 1: The modern quantization pipeline is typically a two-stage process consisting of (1) transformations that make weights and/or activations more amenable to quantization, followed by (2) rounding functions that map weights and/or activations onto a quantization grid.
Before presenting the details of our approach, we first provide a short review of prior works that focus on the two key aspects of quantization we have mentioned: transformation techniques and rounding schemes. Figure 1 illustrates how these two categories interact within the quantization pipeline.
Methods based on transformations. Many recent works propose transformations of weights and/or activations to facilitate quantization. One line of work, initially proposed for MobileNets (Nagel et al., 2019), exploits scaling invariance in neural network compute graphs to equalize the range or precision of weights and activations before quantization. Recent variants leverage scale invariance to redistribute quantization difficulty between weights and activations, with various proposals for learning scales or ranges based on custom objective functions (Xiao et al., 2023; Shao et al., 2024; Lin et al., 2024). Another line of work uses rotations within a compute graph to normalize weight and activation distributions, initially leveraging random orthogonal rotations to promote weight incoherence (Chee et al., 2023). Recent variants employ efficient Hadamard rotations (Tseng et al., 2024a; Ashkboos et al., 2024), Stiefel manifold optimizations (Liu et al., 2025; Hu et al., 2025), and rotation expansion techniques (Adepu et al., 2024; Franco et al., 2025a).
Finally, distinct from these invariance-based approaches, MagR (Zhang et al., 2024) directly minimizes the $\ell_{\infty}$ norm of weights via proximal gradient descent to reduce dynamic range before quantization. While we do not introduce novel transformations of this type in this work, we demonstrate that existing transformations can be combined with our proposed method.
Methods based on rounding. The earliest line of work on rounding relies on continuous optimization strategies based on gradient descent (Nagel et al., 2020). Although more recent methods exist (Hubara et al., 2021; Li et al., 2021), they are not commonly evaluated on LLMs due to computational cost. Early work on LLMs focused on grid scaling or shifting to reduce weight quantization error; for example, LLM.int8() (Dettmers et al., 2022) and ZeroQuant (Yao et al., 2022) directly round to nearest after heuristically selecting the quantization grid, i.e., bit width, scaling factors and zero points. Recent variants have explored more intricate quantization grids, e.g., vector quantization (Tseng et al., 2024a) or trellis codes (Tseng et al., 2024b), that exploit knowledge or assumptions of the data. The most relevant line of work to ours adopts principled discrete optimization using greedy, sequential rounding strategies to select quantized weights to minimize the layer-wise reconstruction error, and includes OBQ (Frantar and Alistarh, 2022), OPTQ (Frantar et al., 2023), GPFQ (Lybrand and Saab, 2021; Zhang et al., 2023), and LDLQ (Chee et al., 2023). Qronos falls within this category.
Notation. Throughout the paper, the weight matrix of a layer is denoted by $W\in\mathbb{R}^{N\times N^{\prime}}$, where each of the $N^{\prime}$ columns represents a $N$-dimensional channel. $\mathcal{A}$ denotes the discrete quantization grid (or alphabet) used for weight quantization, and $\mathcal{Q}$ denotes the corresponding RTN operator associated with $\mathcal{A}$, given by
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{Q}(W):=s\cdot\left(\text{clip}\left(\left\lceil\frac{{W}}{s}\right\rfloor+z;\min\mathcal{A},\max\mathcal{A}\right)-z\right)}.$ (1)
Here, $\textrm{clip}(x;a_{\min},a_{\max})=\min\{\max\{x,a_{\min}\},a_{\max}\}$, while the quantization step size (or scaling factor) is denoted by $s$ and the quantization grid is shifted by an offset denoted by $z$, often referred to as a zero point. In this work, $s,z\in\mathbb{R}^{N^{\prime}}$, and we specify our selection of $s$ and $z$ for the various settings in Section 4. Depending on the setting, different channels (i.e., columns) of the weight matrix may have different quantization grids and RTN operators due to scaling factor granularity (e.g., per-channel or per-group quantization). With a slight abuse of notation, we use the same symbols $\mathcal{A}$ and $\mathcal{Q}$ for all entries. When quantizing $W$, we use $X\in\mathbb{R}^{m\times N}$ to denote the input calibration dataset of $m$ samples (e.g., tokens) for the layer, resulting from the original pre-trained model, and $\widetilde{X}\in\mathbb{R}^{m\times N}$ to denote the input calibration dataset coming from the partially quantized model. Given a vector $v\in\mathbb{R}^{n}$, we use $v_{i}$ for its $i$-th entry, $v_{\geq j}$ for the subvector $(v_{j},\dots,v_{n})^{\top}$, and we define $v_{\leq j}$ analogously. $\|v\|$ is the Euclidean norm of $v$. Given a matrix $A\in\mathbb{R}^{m\times n}$, we use $A_{i}$ to denote its $i$-th column. We use $A_{\geq j}$ to denote the submatrix $(A_{j},\dots,A_{n})$. Similarly, we use $A_{\mathcal{I},\mathcal{J}}$ to represent the submatrix of $A$ consisting of rows indexed by $\mathcal{I}$ and columns indexed by $\mathcal{J}$. For example, $A_{\geq 2,\geq 2}$ will denote the submatrix of $A$ obtained by removing the first row and the first column. $\|A\|_{F}$ denotes the Frobenius norm of $A$. We use $\text{col}(A)$ to denote the column space of $A$. $P_{A}$ is the orthogonal projection onto $\text{col}(A)$, and $P_{A^{\perp}}$ the projection onto its orthogonal complement. Throughout this paper, all indices start from 1, following the standard mathematical convention.
Layer-wise reconstruction and error correction. Data-driven weight quantization methods typically aim to approximately minimize [^2] Equation 2 is an instance of integer least-squares problems, which are known to be NP-hard (Hassibi and Vikalo, 2002). Thus, the best that one can hope for are approximate solutions. the layer-wise reconstruction error given by
$\min_{Q\in\mathcal{A}^{N\times N^{\prime}}}\|XW-XQ\|_{F}^{2}.$ (2)
At an arbitrary layer, the goal is to compute a quantized weight matrix $Q\in\mathcal{A}^{N\times N^{\prime}}$ by approximately solving (2) to preserve the output activations $XW$ under quantization. In practice, however, quantizing weights in earlier layers affects the input to subsequent layers. Let $\widetilde{X}\in\mathbb{R}^{m\times N}$ denote the activation matrix produced by a partially quantized model, where earlier layers have already been quantized. To account for the propagation of quantization error, we use a modified formulation, instead of (2), that targets the mismatch between the original output $XW$ and $\widetilde{X}Q$ by approximately solving
$\min_{Q\in\mathcal{A}^{N\times N^{\prime}}}\|XW-\widetilde{X}Q\|_{F}^{2}.$ (3)
The type of mismatch in this formulation is typically not addressed in the literature but arises naturally in both weight-only and weight-activation quantization settings. For instance, in weight-only quantization, $\widetilde{X}$ arises as the output of previously quantized layers, while in weight-activation quantization, one may encounter $\mathcal{Q}(\widetilde{X})$ rather than $\widetilde{X}$ if activations are quantized. Throughout this paper, we use the notation $(X,\widetilde{X})$ to refer generically to mismatched input pairs.
3 Qronos
We begin by describing the iterations associated with Qronos in Section 3.1. The iterations follow a disciplined and mathematically interpretable framework that alternates between error correction and diffusion using optimal update rules. We then prove that the explicit solutions to these minimization problems admit the efficient implementation detailed in Algorithm 1. In Section 3.2, we provide deeper intuition behind Qronos in the context of previous state-of-the-art rounding algorithms, namely GPFQ and OPTQ. Meanwhile, we also derive a novel interpretation of OPTQ in Appendix F, which shows that it corrects the cumulative quantization error incurred over all the previous iterations.
Algorithm 1 Qronos
$H^{-1}=(\widetilde{X}^{\top}\widetilde{X})^{-1}=LL^{\top}$
for every $w$ in $W$ (in parallel) do
   $q=\mathbf{0}^{N}$
   $w^{(0)}\leftarrow\text{copy}(w)$
   $q_{1}=\mathcal{Q}\left(\dfrac{\widetilde{X}_{1}^{\top}(Xw-\widetilde{X}_{\geq 2}w^{(0)}_{\geq 2})}{\|\widetilde{X}_{1}\|^{2}}\right)$triangleright Calc. $q_{1}$ from (8), use (16) for a memory efficient ver.
   $w^{(1)}_{\geq 2}=\widetilde{X}_{\geq 2}^{\dagger}\left(Xw-\widetilde{X}_{1}q_{1}\right)$triangleright Calc. $w^{(1)}_{\geq 2}$ from (9), see (17) for a memory efficient ver.
   for $t=2$ to $N$ do $\triangleright$ Efficient implementation (Lemma 3.2)
     $q_{t}=\mathcal{Q}(w^{(t-1)}_{t})$
     $w^{(t)}_{\geq t+1}=w^{(t-1)}_{\geq t+1}-L_{\geq t+1,t}\cdot(w^{(t-1)}_{t}-q_{t})/L_{tt}$
   end for
end for
return $Q$
3.1 Algorithm and Efficient Implementation
Let us first note that Qronos can process each column $w\in\mathbb{R}^{N}$ of $W\in\mathbb{R}^{N\times N^{\prime}}$ independently and in parallel to produce each column $q\in\mathcal{A}^{N}$ of $Q\in\mathcal{A}^{N\times N^{\prime}}$. Ideally, the goal is to find $q$ that minimizes $\frac{1}{2}\|Xw-\widetilde{X}q\|^{2}$. Since this problem is NP-hard, we propose an efficient sequential algorithm to approximate its solution. At each iteration, Qronos first selects the quantized weight that optimally corrects the current approximation error, holding the remaining weights fixed; see (4) below. It then updates the unquantized weights to optimally compensate for the rounding error, a process we refer to as error diffusion; see (5).
Let $w$, without superscripts or subscripts, denote the original unquantized weights. After determining $q_{t-1}$, let $w^{(t-1)}_{\geq t}$ represent the updated unquantized weights corresponding to indices $t$ through $N$. The full state of the algorithm after step $t-1$ is thus given by the vector $w^{(t-1)}=(q_{\leq t-1},w^{(t-1)}_{\geq t}),$ with the initialization $w^{(0)}=w$. At step $t$, the algorithm alternates between selecting $q_{t}$ through error correction and updating the remaining weights through error diffusion. The update rules are given by
$q_{t}$ $=\underset{p\in\mathcal{A}}{\mathrm{argmin}}\,\frac{1}{2}\|Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-p\widetilde{X}_{t}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j}\|^{2},$ (4)
$w^{(t)}_{\geq t+1}$ $=\underset{(v_{t+1},\dots,v_{N})\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\,\frac{1}{2}\|Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}v_{j}\widetilde{X}_{j}\|^{2}.$ (5)
These optimization problems admit the following closed-form solutions (see Proposition C.1):
$q_{t}$ $=\mathcal{Q}\left(\frac{\left\langle Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j},\widetilde{X}_{t}\right\rangle}{\|\widetilde{X}_{t}\|^{2}}\right),$ (6)
$w^{(t)}_{\geq t+1}$ $=\widetilde{X}_{\geq t+1}^{\dagger}\left(Xw-\widetilde{X}_{\leq t}q_{\leq t}\right).$ (7)
While these expressions follow directly from the optimization problems, computing $q_{t}$ and $w^{(t)}_{\geq t+1}$ in this form is not computationally efficient and scales poorly, as we will show in Section 4.3. To address this, we present Theorem 3.1, which shows that for all $t\geq 2$, $q_{t}$ can be computed via RTN, enabling a simpler implementation. In Lemma 3.2, we further show that the update for $w^{(t)}_{\geq t+1}$ also admits an efficient implementation using Cholesky decomposition to solve the associated least-squares problem. Together, these results yield the practical and scalable implementation summarized in Algorithm 1.
Theorem 3.1.
Let $(q_{t},w^{(t-1)}_{\geq t})$ be the iterates generated by (4) and (5), with initialization $w^{(0)}_{\geq 1}=w$. Define an alternative sequence $(\hat{q}_{t},\hat{w}^{(t-1)}_{\geq t})$ using the same initialization $\hat{w}^{(0)}_{\geq 1}=w$, by setting
$\hat{q}_{1}$ $=\underset{p\in\mathcal{A}}{\arg\min}\,\frac{1}{2}\|Xw-p\widetilde{X}_{1}-\sum_{j=2}^{N}w_{j}\widetilde{X}_{j}\|^{2},$ (8)
$\hat{w}^{(1)}_{\geq 2}$ $=\underset{(v_{2},\dots,v_{N})\in\mathbb{R}^{N-1}}{\arg\min}\,\frac{1}{2}\|Xw-\hat{q}_{1}\widetilde{X}_{1}-\sum_{j=2}^{N}v_{j}\widetilde{X}_{j}\|^{2},$ (9)
and, for $t=2,\dots,N$, define
$\hat{q}_{t}$ $=\mathcal{Q}(\hat{w}^{(t-1)}_{t}),$ (10)
$\hat{w}^{(t)}_{\geq t+1}$ $=\underset{(v_{t+1},\dots,v_{N})\in\mathbb{R}^{N-t}}{\arg\min}\,\frac{1}{2}\|(\hat{q}_{t}-\hat{w}^{(t-1)}_{t})\widetilde{X}_{t}+\sum_{j=t+1}^{N}(v_{j}-\hat{w}^{(t-1)}_{j})\widetilde{X}_{j}\|^{2}.$ (11)
Then for $t=1,\dots,N$, the two procedures yield identical iterates: $(q_{t},w^{(t-1)}_{\geq t})=(\hat{q}_{t},\hat{w}^{(t-1)}_{\geq t})$.
Starting from the second iteration, Theorem 3.1 shows that the updates in (4) and (5) can be equivalently reformulated as (10) and (11), respectively. This reformulation allows $q_{t}$ to be obtained via RTN for $t\geq 2$, followed by an adjustment of the remaining weights using only the (potentially quantized) activation matrix $\widetilde{X}$ to compensate for the one-step quantization error $(q_{t}-w^{(t-1)}_{t})\widetilde{X}_{t}$.
To further accelerate this adjustment step, we now present Lemma 3.2, which establishes the equivalence of the update in (11) (for $t\geq 2$) with a Cholesky-based least-squares solution [^3] We do not claim that Lemma 3.2 is novel, though we were unable to find it stated explicitly in the literature.. For notational simplicity, we slightly abuse the indexing by treating $t=2$ as a ‘restart.’
Lemma 3.2 (Equivalence of Least-Squares Formulation and Cholesky Formulation).
Assume that $H=X^{\top}X$ is invertible, and let $H^{-1}=LL^{\top}$ denote its Cholesky decomposition, with $L$ lower triangular. Then, starting from $w^{(0)}=w$, the update rules
$q_{t}$ $=\mathcal{Q}(w^{(t-1)}_{t}),$ (12)
$w^{(t)}_{\geq t+1}$ $=\underset{(v_{t+1},\dots,v_{N})\in\mathbb{R}^{N-t}}{\arg\min}\,\frac{1}{2}\|(q_{t}-w^{(t-1)}_{t})X_{t}+\sum_{j=t+1}^{N}(v_{j}-w^{(t-1)}_{j})X_{j}\|^{2}$ (13)
are equivalent to the Cholesky-based iterations
$q_{t}$ $=\mathcal{Q}(w^{(t-1)}_{t}),$ (14)
$w^{(t)}_{\geq t+1}$ $=w^{(t-1)}_{\geq t+1}+\Delta^{(t)},$ (15)
where
$\Delta^{(t)}=-(w^{(t-1)}_{t}-q_{t})\frac{L_{\geq t+1,\,t}}{L_{tt}}\in\mathbb{R}^{N-t}.$
Lemma 3.2 completes our reduction of the original updates (4) and (5) to the equivalent Algorithm 1. A runtime analysis comparing this efficient implementation (Algorithm 1) with a direct evaluation of the closed-form solutions in (6) and (7) is presented in Section 4.3.
At iteration 1 of Algorithm 1, both $q_{1}$ and $w^{(1)}_{\geq 2}$ depend on $\widetilde{X},X\in\mathbb{R}^{m\times N}$, yielding memory requirements of $\mathcal{O}(mN)$, often where $m\gg N$; for example, Llama3.1-8B requires over 30 GB just to store 128 samples of 2048-token sequences at float32, not including weight storage. Observing that OPTQ has memory requirements of $\mathcal{O}(N^{2})$, we can rewrite the first iteration of Algorithm 1 using only square matrices as below, where $G=\widetilde{X}^{T}X\in\mathbb{R}^{N\times N}$ and $H=\widetilde{X}^{T}\widetilde{X}\in\mathbb{R}^{N\times N}$; see Proposition C.2 for a justification.
$q_{1}$ $=\mathcal{Q}\left(\frac{G_{1,\geq 1}w-H_{1,\geq 2}w^{(0)}_{\geq 2}}{H_{11}}\right)$ (16)
$w^{(1)}_{\geq 2}$ $=(H_{\geq 2,\geq 2})^{-1}\left(G_{\geq 2,\geq 1}w-H_{\geq 2,1}q_{1}\right)$ (17)
Note that calculating $G$ and $H$ does not require storing $\widetilde{X},X$, as one can sequentially accumulate the outer products of each of the $m$ samples. Thus, this square matrix formulation reduces peak memory requirements to $\mathcal{O}(N^{2})$, yielding an 18$\times$ reduction in the case of Llama3.1-8B. We note that Colbert et al. (2024) similarly identify a memory optimization for GPFQ, but use singular value decompositions that may not scale well with $N$.
3.2 Theoretical Interpretation and Intuition
Theorem 3.1 and Lemma 3.2 connect the initial disciplined optimization formulation of Qronos, which alternates between (4) and (5), with its efficient practical implementation, described in Algorithm 1. These results guarantee that Qronos is simultaneously interpretable and scalable, explicitly correcting error from the mismatched input pairs $X$ and $\widetilde{X}$. Here, we provide deeper intuition in the context of previous state-of-the-art rounding algorithms, namely GPFQ and OPTQ.
GPFQ is a data-driven error-correcting PTQ method proposed in Lybrand and Saab (2021) and further analyzed in Zhang et al. (2023); Zhang and Saab (2023). When quantizing $w$, it interprets $Xw$ as the endpoint of the path $\sum_{j=1}^{t}w_{j}X_{j}$ for $t=1,...,N$, and handles mismatched inputs by aiming to match $\sum_{j=1}^{t}w_{j}X_{j}$ and $\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}$ for all $t$. More precisely, $q_{t}$ is selected as $\arg\min_{p\in\mathcal{A}}\|\sum_{j=1}^{t}w_{j}X_{j}-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-p\widetilde{X}_{t}\|_{2}^{2}$, as its name (greedy path following quantization) suggests.
Although path following handles the case when $X=\widetilde{X}$ well, additional considerations are required when $X\neq\widetilde{X}$ since, in such a case, the tails of the two paths generally do not align when $\sum_{i=t+1}^{N}w_{i}(X_{i}-\widetilde{X}_{i})\neq 0$. Qronos handles this drawback by adopting a natural remedy to replace the unquantized weights $w_{i},i\geq t+1$ by auxiliary weights $w^{(t)}_{i},i\geq t+1$ so that
$\sum_{i=1}^{t}q_{i}\widetilde{X}_{i}\ +\sum_{i=t+1}^{N}w^{(t)}_{i}\widetilde{X}_{i}\ \approx\ Xw=\sum_{i=1}^{N}w_{i}X_{i}.$
In fact, OPTQ (Frantar et al., 2023) explores a similar weight update idea, but only in the case where $X=\widetilde{X}$, by modifying the remaining unquantized weights after $q_{t}$ is selected. The Cholesky reformulation used in Algorithm 1 also resembles the key mechanism in OPTQ (Algorithm 2, omitting lazy batch updates and dampening, for simplicity). In this way, the runtime of Qronos scales similarly to OPTQ while also explicitly addressing the mismatch between $X$ and $\widetilde{X}$; see Section 4.3 for details. This unexpected connection of Qronos to OPTQ also allows us to derive a novel interpretation of OPTQ in Appendix F, which shows that it corrects the cumulative quantization error incurred over all the previous iterations. We believe this result is of independent interest, due to the widespread use of OPTQ. The proofs for all results are presented in the appendix.
4 Experiments
The core contribution of this work is Qronos—our principled data-driven rounding algorithm that alternates between (1) explicitly correcting quantization error due to both the weights and activations, and (2) diffusing excess error into future weights yet to be quantized. Thus, our primary comparison metric is preserving model quality in challenging quantization scenarios. We design our experiments to isolate the impact of the rounding function (stage 2 in Figure 1), while varying the quantization transforms (stage 1 in Figure 1), as further discussed in Sections 4.1 and 4.2.
Models & Datasets. We conduct experiments on Llama3 (Grattafiori et al., 2024) models using WikiText2 (Merity et al., 2016) for evaluation. We leverage the unmodified implementations made publicly available via Huggingface (Wolf et al., 2020). We use the base foundation model checkpoints for our main results and provide some results with instruction fine-tuned checkpoints in Appendix A. We use LightEval (Fourrier et al., 2023) to evaluate generalization via 5 zero-shot reasoning tasks: ARC (challenge and easy) (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), and Winogrande (Sakaguchi et al., 2021), and report the normalized average accuracy.
Setup. We implement Qronos in PyTorch (Paszke et al., 2019) using the Brevitas quantization library (Franco et al., 2025b), and quantize all models using a single AMD MI210 GPU with 64 GB of memory [^4] AMD, AMD Instinct, and combinations thereof are trademarks of Advanced Micro Devices, Inc.. Unless otherwise specified, we construct our calibration dataset using 128 random sequences of 2048 tokens sampled from the WikiText2 dataset for all data-driven PTQ algorithms. We compare Qronos against RTN and the unmodified Brevitas implementations of OPTQ and GPFQ, also leveraging the unmodified Brevitas implementations of the various quantization transforms. We provide quantization transform hyperparameter details in Appendix B, as well as ablation studies.
Baselines. Our baselines are RTN, OPTQ, and GPFQ. For OPTQ, we use the standard dampened covariance matrix $\widetilde{H}=H+\lambda I$, where $\lambda$ is 1% of the average diagonal of $H$. We similarly use a dampened covariance matrix for Qronos, but choose $\lambda$ to be based on the maximum singular value of $H$ such that $\lambda=\alpha\cdot\sigma_{1}$, which limits the condition number of $\widetilde{H}$ to be less than $\alpha^{-1}$. We select $\alpha=1e^{-6}$ for weight-only quantization and $\alpha=1e^{-3}$ for weight-activation quantization. Additionally, we apply OPTQ, GPFQ, and Qronos block-by-block; for GPFQ and Qronos this corresponds to setting $\widetilde{X}=X$ at the beginning of each block. Note that this corresponds precisely to the original OPTQ proposal (Frantar et al., 2023), and has become common practice for GPFQ (Colbert et al., 2024). Finally, when quantizing weights sequentially, we do so in descending order of the diagonals of $H$, as is also now common practice (IST-DASLab, 2022; Franco et al., 2025b).
Table 2: Weight-only quantization of Llama3 foundation models. We individually apply various quantization transforms (stage 1 in Figure 1) to isolate the impact of different rounding functions (stage 2) when quantizing to 3 and 4 bits, respectively denoted W3 and W4.
    W3 W4
    WikiText2 ($\downarrow$) 0-shot ($\uparrow$) WikiText2 ($\downarrow$) 0-shot ($\uparrow$)
Stage 1 Stage 2     1B 3B 8B     1B 3B 8B     1B 3B 8B     1B 3B 8B
BF16 -     8.9 7.1 5.9     59.4 67.5 74.4     8.9 7.1 5.9     59.4 67.5 74.4
None RTN     2e4 1e4 3e4     32.3 32.4 32.6     18.0 10.1 8.4     49.1 60.8 67.4
OPTQ     42.5 13.8 11.4     37.5 48.1 53.8     10.4 7.8 6.5     54.3 63.4 71.0
GPFQ     35.3 13.4 11.1     35.7 49.9 53.5     10.4 7.8 6.5     56.0 65.2 71.2
Qronos     22.8 11.3 9.3     39.5 53.1 56.7     10.1 7.6 6.4     56.2 64.5 72.0
Smooth Quant RTN     6e3 9e3 5e4     32.7 32.9 31.4     15.2 9.6 8.1     51.4 61.5 67.5
OPTQ     29.6 13.6 12.6     37.0 46.9 47.3     10.4 7.9 6.6     56.2 65.3 70.2
GPFQ     30.1 14.7 12.9     36.5 44.8 45.4     10.8 7.9 6.7     53.9 64.4 69.9
Qronos     19.1 11.6 10.3     40.7 50.6 50.5     10.3 7.8 6.5     56.7 64.8 70.2
MagR RTN     2e3 2e3 5e4     33.8 33.5 35.1     13.8 10.3 7.2     53.1 58.1 69.7
OPTQ     20.1 12.9 8.1     44.2 45.6 59.7     10.3 8.0 6.5     56.4 60.0 69.0
GPFQ     21.0 14.0 8.3     43.9 48.4 61.7     10.4 8.0 6.5     55.4 61.1 70.3
Qronos     16.9 11.8 7.8     46.6 51.2 60.0     10.1 8.0 6.4     56.2 61.1 70.4
HIP RTN     7e2 3e2 1e2     34.2 33.3 36.3     13.8 8.8 7.2     52.0 62.8 70.0
OPTQ     16.1 10.3 8.6     44.1 56.6 58.8     9.9 7.6 6.3     56.8 66.1 72.1
GPFQ     16.6 10.4 8.6     44.9 54.8 58.9     9.9 7.6 6.3     56.5 65.7 72.0
Qronos     12.9 9.3 7.8     48.1 59.6 62.2     9.6 7.5 6.2     57.1 65.9 71.0
4.1 Weight-Only Quantization
We first present state-of-the-art 2-bit and 1.58-bit results for weight-only PTQ on Llama3, controlling for the quantization transform and grid selection while varying the rounding function. We quantize weights using the standard asymmetric weight quantizer (Frantar et al., 2023), where scaling factor $s$ and zero point $z$ are defined per-channel on a scaled min-max grid such that $s=\beta\cdot(\max(w)-\min(w))/(2^{b}-1)$ and $z=\beta\cdot\min(w)/(2^{b}-1)$. Following the analysis of Zhang et al. (2024), we choose $\beta=0.8$ when quantizing to 2 bits or fewer. We combine Hadamard-based incoherence processing (HIP) (Tseng et al., 2024a; Ashkboos et al., 2024) with weight magnitude reduction (MagR) (Zhang et al., 2024) to jointly act as our quantization transform, as they are both known to be effective at few-bit weight quantization (Chee et al., 2023; Adepu et al., 2024). We present our results in Table 1, as well as the BF16 baselines, and highlight that Qronos consistently outperforms existing rounding methods. For example, when compared to OPTQ, Qronos provides a 1.4$\times$ reduction in WikiText2 perplexity and $+$3.3% increase in average zero-shot accuracy for Llama3.2-1B at 2 bits, and a massive improvement in perplexity (4.9$\times$) at 1.58 bits. We provide additional 2-bit and 1.58-bit results with $\beta=1$ in Section B.1.
Next, we present state-of-the-art 3-bit and 4-bit weight-only PTQ results (denoted W3 and W4, respectively) while independently demonstrating compatibility with 3 notable quantization transforms: SmoothQuant (Xiao et al., 2023), MagR, and HIP. Table 2 shows the results across three models in the Llama3 family. For both W3 and W4, we use $\beta=1$. Qronos consistently provides higher quality quantized models than RTN, OPTQ, and GPFQ, as measured in both WikiText2 perplexity and average zero-shot accuracy. Consistent with emerging work on rotation-based quantization transforms (Chee et al., 2023; Tseng et al., 2024a), incoherence processing outperforms other transforms, with HIP + Qronos providing the best overall results. Note that HIP + OPTQ is similar in spirit to QuIP by Theorem 6 in (Chee et al., 2023), which equates LDLQ to OPTQ, with the notable difference that QuIP proposed random orthogonal matrices instead of Hadamard matrices.
4.2 Weight-Activation Quantization
We present 4-bit weight-activation quantization results with and without 4-bit KV cache quantization (denoted W4A4KV16 and W4A4KV4, respectively) while demonstrating compatibility with QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2025). Here, we quantize weights using the standard symmetric weight quantizer with per-channel scaling factors optimized via linear search over the mean square error loss between the full-precision and quantized weights. We quantize activations using the standard asymmetric activation quantizer with dynamic per-token scaling factors and zero points defined on the min-max grid, as is common practice (Liu et al., 2025). When quantizing KV caches, we similarly do so with per-token scaling and zero points. When learning the rotations for SpinQuant, we do so via 100 iterations of Cayley SGD (Li et al., 2020) using 128 random samples from the C4 dataset (Raffel et al., 2020), as further discussed in Appendix B.
Table 3 shows the results across three models in the Llama3 family. Qronos again consistently outperforms RTN, OPTQ, and GPFQ as measured in both WikiText2 perplexity and average zero-shot accuracy. Consistent with emerging work on learned rotations (Liu et al., 2025; Hu et al., 2025; Franco et al., 2025a), SpinQuant outperforms QuaRot, with SpinQuant + Qronos providing the best overall results with and without KV cache quantization. We remark that our experiments use per-token quantization for both the activations and KV caches, while Ashkboos et al. (2024) and Liu et al. (2025) both use per-group scaling for KV cache quantization.
Table 3: Weight-activation quantization of Llama3 models. We individually apply various quantization transforms (stage 1) to isolate the impact of different rounding functions (stage 2).
    W4A4KV16 W4A4KV4
    WikiText2 ($\downarrow$) 0-shot ($\uparrow$) WikiText2 ($\downarrow$) 0-shot ($\uparrow$)
Stage 1 Stage 2     1B 3B 8B     1B 3B 8B     1B 3B 8B     1B 3B 8B
BF16 -     8.9 7.1 5.9     59.4 67.5 74.4     8.9 7.1 5.9     59.4 67.5 74.4
QuaRot RTN     22.0 12.6 9.6     45.4 55.0 62.6     41.8 22.0 15.9     41.5 49.8 57.4
OPTQ     14.3 9.8 8.0     50.4 59.9 66.7     19.8 14.3 10.3     45.8 56.2 64.1
GPFQ     13.6 9.3 7.6     50.9 60.9 67.6     22.0 14.7 11.4     43.3 53.9 59.8
Qronos     13.2 9.1 7.4     50.9 61.5 68.9     17.8 11.6 9.3     47.8 57.3 64.8
SpinQuant RTN     20.1 12.8 9.3     47.4 57.5 63.7     34.7 20.3 13.3     42.6 52.5 60.4
OPTQ     13.6 9.4 7.7     50.5 60.2 67.8     17.7 11.9 10.5     49.4 57.2 63.4
GPFQ     13.5 9.2 7.6     50.8 61.0 66.6     21.5 12.3 10.8     44.1 53.6 60.7
Qronos     12.4 8.7 7.2     52.1 62.1 68.4     16.3 11.1 8.7     48.2 58.2 66.0
(a) Runtime of Rounding Algorithm
(b) Runtime of Quantization Pipeline
Figure 2: We compare the runtime of (a) the rounding algorithm and (b) the overall quantization pipeline as we scale the input features $N$, as measured on an AMD MI210. We average all measurements over 3 seeds and normalize to the runtime of OPTQ where $N=32$.
4.3 Runtime Analysis
Qronos acts as a drop-in replacement at stage 2 of the quantization pipeline shown in Figure 1. We remark that Qronos (and rounding algorithms in general) leave the compute graph unaltered, thus there is no inference overhead on top of that introduced by the quantization transform in stage 1. Since prior works already profile the downstream inference overhead of various compute graph alterations (e.g., Ashkboos et al. (2024) conclude that optimized Hadamard transforms add at most 7% overhead to Llama2 models), we focus our runtime analysis on the quantization pipeline.
We perform our runtime analysis using a single linear layer as our microbenchmark. We use a calibration set of $m=$10,000 random data sampled from a normal $K$-dimensional Gaussian distribution. The linear layer has $K\in[32,1024]$ inputs with $K/4$ outputs. Figure 2 shows how the runtime of OPTQ, GPFQ, and Qronos scale with $K$, where (a) isolates the algorithm runtime and (b) aggregates the quantization pipeline runtime. To highlight the benefits of our equivalent formulation, we implement a base version of Qronos that uses the iterates for $q_{t}$ and $w_{\geq t+1}^{(t)}$ from (6) and (7). Note that via Theorem 3.1 and Lemma 3.2, Algorithm 1 significantly improves the runtime scaling over the base version to match that of OPTQ, with a 13.8$\times$ reduction in algorithm runtime and a 3.6$\times$ reduction in overall runtime when $K=1024$. Compared with OPTQ, which only needs to collect $X$, GPFQ and Qronos require collecting both $\widetilde{X}$ and $X$ at each layer, which requires two forward passes (with and without quantization) and increases the overall quantization pipeline runtime. In practice, we observe that Qronos takes 2$\times$ longer to quantize Llama3.2-1B when compared to OPTQ; however, from Meta (2024), quantization-aware training via QLoRA (Dettmers et al., 2023) took 1,300 GPU hours for Llama3.2-1B, while Qronos took 45 minutes on average.
5 Conclusions
We introduce Qronos—a new state-of-the-art rounding algorithm that alternates between correcting quantization error in both the weights and activations of previous layers and diffusing error into future weights within the current layer. Qronos is based on an interpretable and disciplined optimization framework, and it demonstrably subsumes and surpasses existing data-driven approaches, namely OPTQ and GPFQ. Our implementation exploits several optimizations that together yield orders of magnitude improvements in memory and compute efficiency. Our experiments isolate the impact of the rounding function in the quantization pipeline while varying transformations on a scaled min-max grid. Our results demonstrate that Qronos consistently offers improvements over previous state-of-the-art methods when quantizing weights, activations, and/or KV caches to 4 bits or fewer. That said, our results are intentionally limited to the scaled min-max quantization grid to focus our experiments on transformations and rounding; we believe our results could be further improved by leveraging weight and activation distributions to design quantization grids that are more effective than the scaled min-max grid used in this work, possibly with non-uniform grids via vector quantization.
6 Reproducibility Statement
We integrate Qronos into the open source Brevitas quantization library. Our code and instructions on how to reproduce the various experiments, including hyperparameters and random seeds, can be found here on GitHub. The code points to all datasets and models used for calibration and evaluation.
Acknowledgment
We gratefully acknowledge partial support by National Science Foundation, via the DMS-2410717 grant. From AMD, we would like to thank Nick Fraser and Giuseppe Franco for their constructive reviews and feedback that made this paper better, and Gabor Sines, Michaela Blott, Syed Naim, Yonas Bedasso, and Max Kiehn for their support.
Appendix A Results on Instruction Fine-Tuned Llama3 Models
Our main results evaluate Llama3 foundation models. Here, we demonstrate that Qronos also maintains the quality of instruction fine-tuned Llama3 models better than existing rounding algorithms.
We first present weight-only PTQ results at 3 and 4 bits in Table 4. As in Section 4.1, we asymmetrically quantize weights to the scaled min-max grid with $\beta=1$ for both W3 and W4. We focus our instruction fine-tuned results on evaluating each rounding algorithm with and without Hadamard-based incoherence processing (HIP) as the quantization transform. As in Section 4.1, we find that HIP + Qronos  consistently provides the highest quality quantized models relative to BF16 counterparts, as measured in both WikiText2 perplexity and zero-shot accuracy.
Finally, we present 4-bit weight-activation quantization results with and without 4-bit KV cache quantization in Table 5. We use the same quantization configuration described in Section 4.2, following [Ashkboos et al., 2024, Liu et al., 2025]. As in Section 4.2, we find that SpinQuant outperforms QuaRot, with SpinQuant + Qronos consistently providing the best overall results.
Table 4: Weight-only quantization of instruction fine-tuned Llama3 models. We apply Hadamard-based incoherence processing (HIP) as our quantization transform (stage 1 in Figure 1) to isolate the impact of different rounding functions (stage 2) when quantizing to 3 and 4 bits, respectively denoted W3 and W4. We also evaluate no quantization transform (i.e., "None").
    W3 W4
    WikiText2 ($\downarrow$) 0-shot ($\uparrow$) WikiText2 ($\downarrow$) 0-shot ($\uparrow$)
Stage 1 Stage 2     1B 3B 8B     1B 3B 8B     1B 3B 8B     1B 3B 8B
BF16 -     12.0 9.2 6.7     59.5 66.4 74.1     12.0 9.2 6.7     59.5 66.4 74.1
None RTN     2e4 4e3 3e4     32.6 33.0 32.2     21.4 12.6 9.1     51.0 62.3 67.6
OPTQ     60.0 16.1 12.2     37.4 49.9 58.2     14.3 9.9 7.3     54.5 63.6 71.8
GPFQ     2e2 16.6 12.9     33.8 50.8 55.3     15.4 9.9 7.3     53.3 64.4 71.5
Qronos     43.8 14.3 10.6     37.5 52.1 60.6     13.8 9.8 7.2     55.5 64.8 72.1
HIP RTN     1e3 3e2 1e2     33.4 35.0 36.9     16.6 10.8 8.0     54.6 63.6 70.8
OPTQ     19.1 12.8 9.3     48.0 58.2 59.0     13.2 9.6 7.1     56.6 64.5 72.1
GPFQ     20.4 12.8 9.6     47.6 57.1 61.1     13.2 9.8 7.2     57.0 65.3 71.9
Qronos     16.6 11.6 8.8     49.9 58.4 64.1     12.8 9.6 7.1     57.6 64.8 72.1
Table 5: Weight-activation quantization of instruction fine-tuned Llama3 models. We individually apply QuaRot and SpinQuant as quantization transforms (stage 1 in Figure 1) to isolate the impact of different rounding functions (stage 2) in weight-activation quantization.
    W4A4KV16 W4A4KV4
    WikiText2 ($\downarrow$) 0-shot ($\uparrow$) WikiText2 ($\downarrow$) 0-shot ($\uparrow$)
Stage 1 Stage 2     1B 3B 8B     1B 3B 8B     1B 3B 8B     1B 3B 8B
BF16 -     12.0 9.2 6.7     59.5 66.4 74.1     12.0 9.2 6.7     59.5 66.4 74.1
QuaRot RTN     25.4 17.1 10.9     50.9 55.8 64.0     35.8 27.5 18.3     46.3 51.2 58.7
OPTQ     17.5 12.9 8.8     51.7 60.3 68.6     23.5 18.0 11.1     48.1 54.8 64.9
GPFQ     17.5 12.0 8.8     52.5 62.6 68.1     26.3 20.8 13.6     46.2 54.1 62.1
Qronos     17.5 11.8 8.4     53.1 61.5 69.1     22.0 15.4 10.6     49.8 58.2 66.3
SpinQuant RTN     23.3 17.1 9.9     47.7 54.4 66.9     32.8 25.2 13.6     44.9 51.1 62.2
OPTQ     16.9 11.2 8.5     52.4 61.2 68.9     21.5 15.0 10.1     49.3 56.6 65.4
GPFQ     17.4 12.0 8.8     52.7 62.2 67.7     24.6 19.3 13.2     47.9 55.2 61.8
Qronos     16.3 11.2 8.2     53.9 63.0 69.9     20.3 14.1 10.1     50.0 58.8 67.2
Appendix B Experiment Details for Quantization Transforms
All experiments use WikiText2 as the calibration set, aside from SpinQuant, which uses C4. To pre-process our calibration dataset, we ensure that the <bos> token always appears as the first token in an input sequence as the recent study by Barbero et al. [2025] suggests removing <bos> during inference may greatly reduce performance if models were trained with <bos> always appearing at the first token; their analysis suggests the Llama3 family of models fits this category. Thus, to quantize our models, we first load the pre-trained checkpoint, then pre-process the dataset(s), then apply the quantization pipeline visualized in Figure 1. For Section 4.1, we intentionally select SmoothQuant [Xiao et al., 2023], Hadamard-based incoherence processing (HIP) [Ashkboos et al., 2024, Tseng et al., 2024a], and MagR [Zhang et al., 2024] as they perform fundamentally different transformations. For Section 4.2, we study QuaRot and SpinQuant due to their prevalence. Here, we describe hyperparameters for the data-driven transforms—SmoothQuant, MagR, and SpinQuant.
SmoothQuant. When applying SmoothQuant, we do so before quantizing weights or activations. In practice, SmoothQuant requires the selection of a hyperparameter to control the scaling optimization criteria. We refer to the SmoothQuant hyperparameter as $\gamma$ so as to not clash with our use of $\alpha$ in Section 4; note that $\gamma\in[0,1]$. In Table 6, we provide the results of a uniform grid search over $\gamma$ when quantizing Llama3.2-1B-Instruct to 4 bits using round-to-nearest (RTN). These results motivate our decision to use $\gamma=0.3$ in all our weight-only PTQ experiments that apply SmoothQuant.
Table 6: Impact of SmoothQuant’s $\gamma$ on Llama3.2-1B-Instruct. We evaluate the impact of the smoothing parameter $\gamma$ on both WikiText2 perplexity and normalized average zero-shot accuracy when quantizing Llama3.2-1B-Instruct to 4 bits using round-to-nearest (RTN).
$\gamma$ 0.2 0.3 0.4 0.5 0.6 0.7 0.8
WikiText2 ($\downarrow$) 24.6 18.6 18.9 21.4 87.0 4e2 3e4
0-shot ($\uparrow$) 50.8 53.3 52.8 52.5 42.8 36.6 32.3
MagR. When applying MagR, we also do so before quantizing weights and activations. When coupled with HIP, we do so after inserting rotations into the compute graph. In practice, MagR requires tuning the $\ell_{\infty}$ penalty; we refer to this hyperparameter as $\theta$, again so as to not clash with our use of $\alpha$ in Section 4. Zhang et al. [2024] tune $\theta$ to Llama2 models, settling on $\theta=0.001$ for their experiments. In Table 7, we provide new results for Llama3.2-1B-Instruct. These results motivate our decision to use $\theta=0.01$ in all our weight-only PTQ experiments that apply MagR.
Table 7: Impact of MagR’s $\theta$ on Llama3.2-1B-Instruct. We evaluate the impact of the penalty parameter $\theta$ on both WikiText2 perplexity and normalized average zero-shot accuracy when quantizing Llama3.2-1B-Instruct to 4 bits using round-to-nearest (RTN).
$\theta$ 0.1 0.01 0.001 0.0001
WikiText2 ($\downarrow$) 74.5 25.4 105.0 216.0
0-shot ($\uparrow$) 44.2 53.0 44.7 42.6
SpinQuant. When applying SpinQuant, Liu et al. [2025] do so after activation (and KV cache) quantization but before weight quantization using an 800-sample calibration dataset; their ablation study demonstrates negligible degradation when using 128 samples. Thus, we employ Cayley SGD on a network where only activations are quantized to optimize the learnable rotations for 100 iterations using a calibration dataset constructed of 128 random samples from the C4 dataset.
B.1 Grid scaling ablation study for 2 bits and fewer
In Section 4.1, we present weight-only PTQ results when quantizing to 2 bits or fewer on the scaled min-max grid. Here, in Table 8, we provide additional results that demonstrate Qronos outperforms other rounding algorithms on another choice of $\beta$. Recall that we jointly apply Hadamard-based incoherence processing (HIP) and weight magnitude reduction (MagR) as quantization transforms before each rounding algorithm. Our results highlight that $\beta=0.8$ (see Table 1) is an overall better choice for scaling the min-max grid in this setting, which is consistent with Zhang et al. [2024], and that Qronos provides the best results on both grids at all bit widths and model sizes.
Table 8: Weight-only quantization of Llama3 models to 2 bits or fewer with $\beta=1$. We jointly apply HIP and MagR as quantization transforms (stage 1 in Figure 1) and compare different rounding functions (stage 2) on the scaled min-max grid (see Section 4). Note that these results complement Table 1, which presents results with $\beta=0.8$.
WikiText2 ($\downarrow$) 0-shot ($\uparrow$)
1B 3B 8B     1B 3B 8B
BF16 - 8.9 7.1 5.9     59.4 67.5 74.4
2-bit RTN 1e4 1e4 2e4     32.4 32.4 32.9
OPTQ 45.3 20.8 18.9     35.2 39.3 41.2
GPFQ 47.5 22.4 17.8     33.9 38.4 39.2
Qronos 24.6 14.9 12.4     38.4 43.4 45.6
1.58-bit RTN 2e5 3e5 6e5     32.0 32.6 32.1
OPTQ 5e3 4e2 3e2     32.5 32.4 32.2
GPFQ 6e2 7e2 5e2     31.2 32.5 32.7
Qronos 79.5 48.3 34.8     32.9 32.8 34.3
Appendix C Preliminary Propositions
Proposition C.1.
The update rule given by
$q_{t}$ $=\underset{p\in\mathcal{A}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-p\widetilde{X}_{t}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j}\|^{2},$
$w^{(t)}_{\geq t+1}$ $=\underset{(v_{t+1},...,v_{N})\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}v_{j}\widetilde{X}_{j}\|^{2}.$
has closed-form expressions
$q_{t}=\mathcal{Q}\left(\frac{\langle Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j},\widetilde{X}_{t}\rangle}{\|\widetilde{X}_{t}\|^{2}}\right)$
and
$w^{(t)}_{\geq t+1}=\widetilde{X}_{\geq t+1}^{\dagger}\left(Xw-\widetilde{X}_{\leq t}q_{\leq t}\right).$
Proof.
For $q_{t}$, the corresponding optimization objective function is a one-dimensional quadratic function of $p$. Since minimizing a quadratic function over a discrete set $\mathcal{A}$ reduces to rounding its real-valued minimizer, we compute the real-valued minimizer
$\frac{\langle Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j},\widetilde{X}_{t}\rangle}{\|\widetilde{X}_{t}\|^{2}}.$
Thus, we obtain the closed-form expression of $q_{t}$,
$q_{t}=\mathcal{Q}\left(\frac{\langle Xw-\sum_{j=1}^{t-1}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}\widetilde{X}_{j},\widetilde{X}_{t}\rangle}{\|\widetilde{X}_{t}\|^{2}}\right),$
where $\mathcal{Q}$ is the round-to-nearest operator.
For $w^{(t)}_{\geq t+1}$, the corresponding optimization problem is an unconstrained least-square problem in the form of $\min_{v\in\mathbb{R}^{N-t}}\|Ax-b\|^{2}$, with $A=\widetilde{X}_{\geq t+1}$ and $b=Xw-\widetilde{X}_{\leq t}q_{\leq t}$. Thus, the minimizer is given by $A^{\dagger}b$, which gives the desired closed-form expression. ∎
Proposition C.2.
The update rule given by
$q_{1}=\mathcal{Q}\left(\dfrac{\widetilde{X}_{1}^{\top}(Xw-\widetilde{X}_{\geq 2}w^{(0)}_{\geq 2})}{\|\widetilde{X}_{1}\|^{2}}\right),$
$w^{(1)}_{\geq 2}=\widetilde{X}_{\geq 2}^{\dagger}\left(Xw-\widetilde{X}_{1}q_{1}\right)$
is equivalent to
$q_{1}$ $=\mathcal{Q}\left(\frac{G_{1,\geq 1}w-H_{1,\geq 2}w^{(0)}_{\geq 2}}{H_{11}}\right)$
$w^{(1)}_{\geq 2}$ $=(H_{\geq 2,\geq 2})^{-1}\left(G_{\geq 2,\geq 1}w-H_{\geq 2,1}q_{1}\right),$
where $G=\widetilde{X}^{T}X\in\mathbb{R}^{N\times N}$ and $H=\widetilde{X}^{T}\widetilde{X}\in\mathbb{R}^{N\times N}$.
Proof.
For $q_{1}$, we have $\widetilde{X}_{1}^{\top}X=(\widetilde{X}^{\top}X)_{1,\geq 1}=G_{1,\geq 1}$. Also, $\widetilde{X}_{1}^{\top}\widetilde{X}_{\geq 2}=(\widetilde{X}^{\top}\widetilde{X})_{1,\geq 2}=H_{1,\geq 2}$. Thus, $\widetilde{X}_{1}^{\top}(Xw-\widetilde{X}_{\geq 2}w^{(0)}_{\geq 2})=G_{1,\geq 1}w-H_{1,\geq 2}w^{(0)}_{\geq 2}$. Further, $\|\widetilde{X}_{1}\|^{2}=(\widetilde{X}^{\top}\widetilde{X})_{11}=H_{11}$. This gives the equivalence for updating $q_{1}$.
For $w^{(1)}_{\geq 2}$, $\widetilde{X}_{\geq 2}$ is given by $(\widetilde{X}_{\geq 2}^{\top}\widetilde{X}_{\geq 2})^{-1}\widetilde{X}_{\geq 2}^{\top}=(H_{\geq 2,\geq 2})^{-1}\widetilde{X}_{\geq 2}^{\top}$. Then
$\widetilde{X}_{\geq 2}^{\dagger}\left(Xw-\widetilde{X}_{1}q_{1}\right)$ $=(H_{\geq 2,\geq 2})^{-1}\widetilde{X}_{\geq 2}^{\top}\left(Xw-\widetilde{X}_{1}q_{1}\right)$
$=(H_{\geq 2,\geq 2})^{-1}\left((\widetilde{X}^{\top}X)_{\geq 2,\geq 1}w-(\widetilde{X}^{\top}\widetilde{X})_{\geq 2,1}q_{1}\right)$
$=(H_{\geq 2,\geq 2})^{-1}\left(G_{\geq 2,\geq 1}w-H_{\geq 2,1}q_{1}\right).$
This gives the equivalence for updating $w^{(1)}_{\geq 2}.$
Appendix D Proof of Theorem 3.1
Proof.
We use induction to prove the theorem. Since at $t=1$ equations (4), (5) and equations (8), (9) are identical, the base case is trivially true. Now we proceed with the induction, assuming $\hat{w}^{(t)}_{\geq t+1}=w^{(t)}_{\geq t+1}$ and $\hat{q}_{t}=q_{t}$.
Using definition (4) and Proposition C.1, we can obtain the closed-form expression,
$q_{t+1}=\mathcal{Q}\left(\frac{\langle Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+2}^{N}w^{(t)}_{j}\widetilde{X}_{j},\widetilde{X}_{t+1}\rangle}{\|\widetilde{X}_{t+1}\|^{2}}\right),$
where $\mathcal{Q}$ is the RTN operator. Next we note that (5), which is used to compute $w^{(t)}_{\geq t+1}$, implies that $Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j}$ is orthogonal to the column space of $\widetilde{X}_{\geq t+1}$. This in turn implies that $\langle Xw-\sum_{j=1}^{t}q_{j}\tilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j},\widetilde{X}_{t+1}\rangle=0$. Then we can compute,
$q_{t+1}$ $=\mathcal{Q}\left(\frac{\langle Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+2}^{N}w^{(t)}_{j}\widetilde{X}_{j},\widetilde{X}_{t+1}\rangle}{\|\widetilde{X}_{t+1}\|^{2}}\right)$
$=\mathcal{Q}\left(\frac{\langle Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j}+w^{(t)}_{t+1}\widetilde{X}_{t+1},\widetilde{X}_{t+1}\rangle}{\|\widetilde{X}_{t+1}\|^{2}}\right)$
$=\mathcal{Q}\left(\frac{\langle w^{(t)}_{t+1}\widetilde{X}_{t+1},\widetilde{X}_{t+1}\rangle}{\|\widetilde{X}_{t+1}\|^{2}}\right)$
$=\mathcal{Q}\left(w^{(t)}_{t+1}\right)=\mathcal{Q}\left(\hat{w}^{(t)}_{t+1}\right)=\hat{q}_{t+1},$
where in the last two inequalities, we used the induction hypothesis $\hat{w}^{(t)}_{\geq t+1}=w^{(t)}_{\geq t+1}$ and the update rule (10).
Next, we prove $\hat{w}^{(t+1)}_{\geq t+2}=w^{(t+1)}_{\geq t+2}.$ We first compute
$w^{(t+1)}_{\geq t+2}$ $=\underset{v_{\geq t+2}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t+1}q_{j}\widetilde{X}_{j}-\sum_{j=t+2}^{N}v_{j}\widetilde{X}_{j}\|^{2}$
$=\underset{v_{\geq t+2}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j}+(w^{(t)}_{t+1}-q_{t+1})\widetilde{X}_{t+1}+\sum_{j=t+2}^{N}(w^{(t)}_{j}-v_{j})\widetilde{X}_{j}\|^{2}.$
Due to the update rule (5), $Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j}$ is orthogonal to the column span of $\widetilde{X}_{\geq t+1}$, hence to $(w^{(t)}_{t+1}-q_{t+1})\widetilde{X}_{t+1}+\sum_{j=t+2}^{N}(w^{(t)}_{j}-v_{j})\widetilde{X}_{j}$. Then, we have
$w^{(t+1)}_{\geq t+2}$ $=\underset{v_{\geq t+2}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}q_{j}\widetilde{X}_{j}-\sum_{j=t+1}^{N}w^{(t)}_{j}\widetilde{X}_{j}+(w^{(t)}_{t+1}-q_{t+1})\widetilde{X}_{t+1}+\sum_{j=t+2}^{N}(w^{(t)}_{j}-v_{j})\widetilde{X}_{j}\|^{2}$
$=\underset{v_{\geq t+2}}{\mathrm{argmin}}\frac{1}{2}\|(\hat{w}^{(t)}_{t+1}-\hat{q}_{t+1})\widetilde{X}_{t+1}+\sum_{j=t+2}^{N}(\hat{w}^{(t)}_{j}-v_{j})\widetilde{X}_{j}\|^{2}=\hat{w}^{(t+1)}_{\geq t+2},$
where we used the Pythagorean theorem, the induction hypothesis $\hat{w}^{(t)}_{\geq t+1}=w^{(t)}_{\geq t+1}$, and the fact $q_{t+1}=\hat{q}_{t+1}$. This completes the induction. ∎
Appendix E Proof of Lemma 3.2
Throughout this section, we denote $H_{\geq t,\geq t}=X_{\geq t}^{\top}X_{\geq t}\in\mathbb{R}^{(N-t+1)\times(N-t+1)}$ and $H^{-1}_{\geq t,\geq t}=(X_{\geq t}^{\top}X_{\geq t})^{-1}\in\mathbb{R}^{(N-t+1)\times(N-t+1)}$. We will begin with a few preliminary lemmas before we prove Lemma 3.2. While some of these lemmas may already be known, we are not aware of any rigorous proofs in the literature. Thus, we provide our proofs here for completeness.
Lemma E.1.
Denote by $[H^{-1}_{\geq t,\geq t}]_{11}$ the first entry of $H^{-1}_{\geq t,\geq t}$ and by $[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}\in\mathbb{R}^{N-t}$ the first column of $H^{-1}_{\geq t,\geq t}$ albeit with the first entry removed. Then
$(X_{\geq t+1}^{\top}X_{\geq t+1})^{-1}X_{\geq t+1}^{\top}X_{t}=-\frac{[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}}{[H^{-1}_{\geq t,\geq t}]_{11}}.$
Proof.
[…]
$(X_{\geq t+1}^{\top}X_{\geq t+1})^{-1}X_{\geq t+1}^{\top}X_{t}=-\frac{\mathbf{b}}{r}$
and finishes the proof. ∎
The next lemma establishes how one can efficiently compute $H^{-1}_{\geq t+1,\geq t+1}$ from $H^{-1}_{\geq t,\geq t}$.
Lemma E.2.
$H^{-1}_{\geq t+1,\geq t+1}$ can be efficiently computed from $H^{-1}_{\geq t,\geq t}$ via
$H^{-1}_{\geq t+1,\geq t+1}=\left(H^{-1}_{\geq t,\geq t}-\frac{1}{[H^{-1}_{\geq t,\geq t}]_{11}}[H^{-1}_{\geq t,\geq t}]_{\geq 1,1}[H^{-1}_{\geq t,\geq t}]_{1,\geq 1}\right)_{\geq 2,\geq 2}.$
We note that this is a simple rank-$1$ update followed by a submatrix slicing.
Proof.
We first recall a more general inverse formula for $2\times 2$ block matrix using the Schur complement. Consider the $2\times 2$ block matrix
$M=\begin{pmatrix}A&B\\ C&D\end{pmatrix}.$
When $A$ is invertible, the inverse of $M$ is given by
$M^{-1}=\begin{pmatrix}A^{-1}+A^{-1}BS^{-1}CA^{-1}&-A^{-1}BS^{-1}\\ -S^{-1}CA^{-1}&S^{-1}\end{pmatrix},$ (18)
where $S=D-CA^{-1}B$ is the Schur complement of $A$ in $M$.
When $A$ is a scalar $a$ and ${M}$ is symmetric, i.e.
${M}=\begin{pmatrix}a&b^{\top}\\ b&D\end{pmatrix},$
this formula becomes
${M}^{-1}=\begin{pmatrix}a^{-1}+a^{-2}b^{\top}S^{-1}b&-a^{-1}b^{\top}S^{-1}\\ -a^{-1}S^{-1}b&S^{-1}\end{pmatrix},$
where $S=D-a^{-1}bb^{\top}$.
By the Sherman–Morrison formula [Horn and Johnson, 2012], we have
$D^{-1}$ $=S^{-1}-\frac{S^{-1}bb^{\top}S^{-1}}{a+b^{\top}S^{-1}b}$
$=S^{-1}-\frac{a^{-2}S^{-1}bb^{\top}S^{-1}}{a^{-1}+a^{-2}b^{\top}S^{-1}b}.$
Returning to our setting where ${M}^{-1}=H^{-1}_{\geq t,\geq t}$ and ${D}^{-1}=H^{-1}_{\geq t+1,\geq t+1}$, we have
$H^{-1}_{\geq t+1,\geq t+1}$ $=[H^{-1}_{\geq t,\geq t}]_{\geq 2,\geq 2}-\frac{1}{[H^{-1}_{\geq t,\geq t}]_{11}}[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}[H^{-1}_{\geq t,\geq t}]_{1,\geq 2}$
$=[H^{-1}_{\geq t,\geq t}]_{\geq 2,\geq 2}-\frac{1}{[H^{-1}_{\geq t,\geq t}]_{11}}\left([H^{-1}_{\geq t,\geq t}]_{\geq 1,1}[H^{-1}_{\geq t,\geq t}]_{1,\geq 1}\right)_{\geq 2,\geq 2}$
$=\left(H^{-1}_{\geq t,\geq t}-\frac{1}{[H^{-1}_{\geq t,\geq t}]_{11}}[H^{-1}_{\geq t,\geq t}]_{\geq 1,1}[H^{-1}_{\geq t,\geq t}]_{1,\geq 1}\right)_{\geq 2,\geq 2}.$
Using the above lemma and Cholesky decomposition [Horn and Johnson, 2012], we can further simplify the right hand side in Lemma E.1 via the following lemma.
Lemma E.3.
Let $H^{-1}=(X^{\top}X)^{-1}$ and $H^{-1}=LL^{\top}$ be its Cholesky decomposition where $L$ is a lower triangular matrix, then
$\frac{[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}}{[H^{-1}_{\geq t,\geq t}]_{11}}=\frac{L_{\geq t+1,t}}{L_{tt}}\in\mathbb{R}^{N-t}$
holds for all $t\in[N-1]$.
Proof.
We first prove that given the Cholesky decomposition $H^{-1}=LL^{\top}$, the Cholesky decomposition of $H^{-1}_{\geq t,\geq t}$ is $H^{-1}_{\geq t,\geq t}=(L_{\geq t,\geq t})(L_{\geq t,\geq t})^{\top}$ for all $t\in[N]$, where $H^{-1}_{\geq t,\geq t}=(X_{\geq t}^{\top}X_{\geq t})^{-1}\in\mathbb{R}^{(N-t+1)\times(N-t+1)}$.
Let us proceed by induction. The base-case when $t=1$ holds by assumption, and we now assume the result holds for $t$. By Lemma E.2, the updated inverse Hessian $H^{-1}_{\geq t+1,\geq t+1}=\left(H^{-1}_{\geq t,\geq t}-\frac{1}{[H^{-1}_{\geq t,\geq t}]_{11}}[H^{-1}_{\geq t,\geq t}]_{\geq 1,1}[H^{-1}_{\geq t,\geq t}]_{1,\geq 1}\right)_{\geq 2,\geq 2}$. Thus,
$\Big{(}(L_{\geq t,\geq t})$ $(L_{\geq t,\geq t})^{\top}-\frac{1}{L_{tt}^{2}}((L_{\geq t,\geq t})_{11}\cdot[L_{\geq t,\geq t}]_{\geq 1,1})((L_{\geq t,\geq t})_{11}\cdot[L_{\geq t,\geq t}]_{\geq 1,1})^{\top}\Big{)}_{\geq 2,\geq 2}$
$=$ $\left((L_{\geq t,\geq t})(L_{\geq t,\geq t})^{\top}-[L_{\geq t,\geq t}]_{\geq 1,1}[L_{\geq t,\geq t}]_{\geq 1,1}^{\top}\right)_{\geq 2,\geq 2}$
$=$ $((L_{\geq t,\geq t})_{\geq 2,\geq 2})((L_{\geq t,\geq t})_{\geq 2,\geq 2})^{\top}$
$=$ $(L_{\geq t+1,\geq t+1})(L_{\geq t+1,\geq t+1})^{\top}$
This finishes the induction and we have Cholesky decomposition $H^{-1}_{\geq t,\geq t}=(L_{\geq t,\geq t})(L_{\geq t,\geq t})^{\top}$ for all $t\in[N]$. To finish the proof, let $M=RR^{\top}$ be the Cholesky decomposition of any positive definite matrix $M$. By a direct computation, the first column of $M$ is $R[R^{\top}]_{\geq 1,1}=R_{11}\cdot[R]_{\geq 1,1}$ and the first entry $M_{11}=R_{11}^{2}$. Then we have $\frac{M_{\geq 1,1}}{M_{11}}=\frac{[R]_{\geq 1,1}}{R_{11}}$ which implies that $\frac{M_{\geq 2,1}}{M_{11}}=\frac{[R]_{\geq 2,1}}{R_{11}}$. In our case, we have $H^{-1}_{\geq t,\geq t}=(L_{\geq t,\geq t})(L_{\geq t,\geq t})^{\top}$ in the place of $M=RR^{\top}$. Thus,
$\frac{[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}}{[H^{-1}_{\geq t,\geq t}]_{11}}=\frac{[L_{\geq t,\geq t}]_{\geq 2,1}}{[L_{\geq t,\geq t}]_{11}}=\frac{L_{\geq t+1,t}}{L_{tt}}.$
With the above preliminary lemmas, now we are ready to prove Lemma 3.2
Proof of Lemma 3.2.
Since we initialize with $w^{(0)}=w$, $q_{1}=\mathcal{Q}(w_{1})$ always holds. Thus the two iterations produce the same $q_{1}$ and $w^{(0)}_{\geq 1}$ We proceed by induction. Assume at step $t$ that $q_{t}$ and $w^{(t-1)}_{\geq t}$ resulting from the update rules (12) and (13) match those following update rules (14) and (15). In order to complete the induction, it suffices to show that (13) and (15) produce the same $w^{(t)}_{\geq t+1}$, which naturally results in the same $q_{t+1}=\mathcal{Q}(w^{(t)}_{t+1})$.
To that end, we note that the optimization problem defined by (13) has a unique least-square solution as $X_{\geq t+1}$ has full column rank. The minimizer is given by
${w}_{\geq t+1}^{(t)}$ $=w_{\geq t+1}^{(t-1)}+(w^{(t-1)}_{t}-q_{t})X_{t+1:}^{\dagger}X_{t}$
$=w_{\geq t+1}+(w^{(t-1)}_{t}-q_{t})(X_{\geq t+1}^{\top}X_{\geq t+1})^{-1}X_{\geq t+1}^{\top}X_{t}$
By Lemma E.1, we have
$(X_{\geq t+1}^{\top}X_{\geq t+1})^{-1}X_{\geq t+1}^{\top}X_{t}=-\frac{[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}}{[H^{-1}_{\geq t,\geq t}]_{11}}.$
Lastly, Lemma E.3 gives us
$\frac{[H^{-1}_{\geq t,\geq t}]_{\geq 2,1}}{[H^{-1}_{\geq t,\geq t}]_{11}}=\frac{L_{\geq t+1,t}}{L_{tt}}\in\mathbb{R}^{N-t}.$
This matches $\Delta_{t+1}$ in (15) and completes our induction. ∎
Appendix F A Novel Interpretation of OPTQ
Algorithm 2 OPTQ: Quantize a layer $W$ given inverse Hessian $H^{-1}=(X^{\top}X)^{-1}$.
1:for every $w$ in $W$ in parallel do
2:   $q=\mathbf{0}^{N}$$\triangleright$ Initialize quantized neuron
3:   $H^{-1}=LL^{\top}$triangleright Perform Cholesky decomposition
4:   for $t=1$ to $N$ do $\triangleright$ Iterate over rows
5:     $q_{t}=\mathcal{Q}(w_{t})$
6:     $w_{\geq t}\leftarrow w_{\geq t}-L_{\geq t,t}\cdot(w_{t}-q_{t})/L_{tt}$$\triangleright$ Update remaining weights
7:   end for
8:end for
9:return $Q$
For our final result of this paper, we observe that updates of $w^{(t)}_{\geq t+1}$ via (13) can be interpreted by observing that the term $(q_{t}-w^{(t-1)}_{t})X_{t}$ represents the error introduced by quantizing $w^{(t-1)}_{t}$. The optimization problem (13) seeks to mitigate this error by adjusting future weights so as to minimize the resulting distortion, measured in the $\ell_{2}$-norm. Notably, this step does not explicitly attempt to correct errors introduced by earlier quantization steps $11,\dots,t-$ 1. However, by combining the proof of Theorem 3.1 in the case when $X=\widetilde{X}$ with Lemma 3.2, we arrive at Corollary F.1, which provides a novel interpretation of OPTQ. It shows—perhaps unexpectedly—that Algorithm 2 optimally corrects the cumulative quantization error incurred over the first $t$ entries of $w$.
Corollary F.1.
The OPTQ update rules from Algorithm 2 are equivalent to
$q_{t}$ $=\underset{p\in\mathcal{A}}{\arg\min}\,\frac{1}{2}\|Xw-\sum_{j=1}^{t-1}q_{j}X_{j}-pX_{t}-\sum_{j=t+1}^{N}w^{(t-1)}_{j}X_{j}\|^{2},$ (19)
$w^{(t)}_{\geq t+1}$ $=\underset{(v_{t+1},\dots,v_{N})\in\mathbb{R}^{N-t}}{\arg\min}\,\frac{1}{2}\|Xw-\sum_{j=1}^{t}q_{j}X_{j}-\sum_{j=t+1}^{N}v_{j}X_{j}\|^{2},$ (20)
with $w^{(0)}_{\geq 1}=w$. In other words, the updated weights and quantized weights at every iteration $t$ that are produced by Algorithm 2 are identical to those produced by equations (19) and (20). In particular, (20) shows that, at each step, the updated OPTQ weights $w^{(t)}_{\geq t+1}$indeed optimally correct for the errors produced by the hitherto quantized sequence $q_{1},...,q_{t}$.
Proof.
The proof is based on induction on both arguments of the trajectory. Let $\{(\hat{w}^{(t-1)}_{\geq t},\hat{q}_{t})\}_{t=1}^{N}$ denote the trajectory generated by update rules (19), (20). And let $\{(w^{(t-1)}_{\geq t},q_{t})\}_{t=1}^{N}$ be the trajectory generated by Algorithm 2. Our goal is to prove $(\hat{w}^{(t-1)}_{\geq t},\hat{q}_{t})=(w^{(t-1)}_{\geq t},q_{t})$ for $t=1,\dots,N$.
By Lemma 3.2, the trajectory $\{(w^{(t-1)}_{\geq t},q_{t})\}_{t=1}^{N}$ generated using Cholesky decomposition in Algorithm 2 can be equivalently regarded as generated from (12), (13). Thus, we will use (12), (13) as the update rule of $w^{(t-1)}_{\geq t}$ and $q_{t}$ in the rest of our proof. In the base case, $\hat{w}^{(0)}_{\geq 1}=w^{(0)}_{\geq 1}$ are both initialized with $w$ and
$\hat{q}_{1}=\underset{p\in\mathcal{A}}{\mathrm{argmin}}\frac{1}{2}\|Xw-pX_{1}-\sum_{j=2}^{N}w^{(0)}_{j}X_{j}\|^{2}=\underset{p\in\mathcal{A}}{\mathrm{argmin}}\frac{1}{2}\|(w_{1}-p)X_{1}\|^{2}=\mathcal{Q}(w_{1})=q_{1}.$
Thus $(w^{(0)}_{\geq 1},q_{1})=(\hat{w}^{(0)}_{\geq 1},\hat{q}_{1})$. Assume $(\hat{w}^{(t-1)}_{\geq t},\hat{q}_{t})=(w^{(t-1)}_{\geq t},q_{t})$ holds true. Now we proceed to prove $(\hat{w}^{(t)}_{\geq t+1},\hat{q}_{t+1})=(w^{(t)}_{\geq t+1},q_{t+1})$.
Step 1: We first prove $\hat{w}^{(t)}_{\geq t+1}={w}^{(t)}_{\geq t+1}$. By construction,
$\hat{w}^{(t)}_{\geq t+1}=\underset{v_{\geq t+1}\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}v_{j}X_{j}\|^{2}.$
For an arbitrary $v_{\geq t+1}\in\mathbb{R}^{N-t}$,
$Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}$ $-\sum_{j=t+1}^{N}v_{j}X_{j}=$
$\underbrace{(Xw-\sum_{j=1}^{t-1}\hat{q}_{j}X_{j}-\sum_{j=t}^{N}\hat{w}^{(t-1)}_{j}X_{j})}_{\textbf{(I)}}+\underbrace{\left((\hat{w}^{(t-1)}_{t}-\hat{q}_{t})X_{t}+\sum_{j=t+1}^{N}(\hat{w}^{(t-1)}_{j}-v_{j})X_{j}\right)}_{\textbf{(II)}}.$
[…]
$\left\|Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}v_{j}X_{j}\right\|^{2}=\left\|\textbf{(I)}\right\|^{2}+\left\|\textbf{(II)}\right\|^{2}.$
Notice that (I) does not depend on $v_{\geq t+1}$. Furthermore, $\hat{w}^{(t-1)}_{\geq t+1}$ and $\hat{q}_{t}$ in (II) can be replaced by ${w}^{(t-1)}_{\geq t+1}$ and $q_{t}$ respectively using our induction hypothesis. Thus,
$\hat{w}^{(t)}_{\geq t+1}$ $=\underset{v_{\geq t+1}\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}v_{j}X_{j}\|^{2}$
$=\underset{v_{\geq t+1}\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|(\hat{w}^{(t-1)}_{t}-\hat{q}_{t})X_{t}+\sum_{j=t+1}^{N}(\hat{w}^{(t-1)}_{j}-v_{j})X_{j}\|^{2}$
$=\underset{v_{\geq t+1}\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|({w}^{(t-1)}_{t}-{q}_{t})X_{t}+\sum_{j=t+1}^{N}({w}^{(t-1)}_{j}-v_{j})X_{j}\|^{2}$
$={w}^{(t)}_{\geq t+1}.$
Step 2: Now we prove $\hat{q}_{t+1}=q_{t+1}$. We just constructed
$\hat{w}^{(t)}_{\geq t+1}=\underset{v_{\geq t+1}\in\mathbb{R}^{N-t}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}v_{j}X_{j}\|^{2}.$
This implies
$Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}\hat{w}^{(t)}_{j}X_{j}=P_{X_{\geq t+1}^{\perp}}(Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j})\in X_{\geq t+1}^{\perp}.$ (21)
By construction, we have
$\hat{q}_{t+1}$ $=\underset{q\in\mathcal{A}}{\mathrm{argmin}}\frac{1}{2}\|Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-qX_{t+1}-\sum_{j=t+2}^{N}\hat{w}^{(t)}_{j}X_{j}\|^{2}$
$=\mathcal{Q}\left(\frac{\langle X_{t+1},Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+2}^{N}\hat{w}^{(t)}_{j}X_{j}\rangle}{\|X_{t+1}\|^{2}}\right).$
Then we can use (21) to deduce
$\frac{\langle X_{t+1},Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+2}^{N}\hat{w}^{(t)}_{j}X_{j}\rangle}{\|X_{t+1}\|^{2}}$
$=\frac{\langle X_{t+1},Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}\hat{w}^{(t)}_{j}X_{j}+X_{t+1}\hat{w}^{(t)}_{t+1}\rangle}{\|X_{t+1}\|^{2}}$
$=\frac{\langle X_{t+1},Xw-\sum_{j=1}^{t}\hat{q}_{j}X_{j}-\sum_{j=t+1}^{N}\hat{w}^{(t)}_{j}X_{j}\rangle}{\|X_{t+1}\|^{2}}+\frac{\langle X_{t+1},X_{t+1}\hat{w}^{(t)}_{t+1}\rangle}{\|X_{t+1}\|^{2}}$
$=\frac{\langle X_{t+1},X_{t+1}\hat{w}^{(t)}_{t+1}\rangle}{\|X_{t+1}\|^{2}}$
$=\hat{w}^{(t)}_{t+1}$
$=w^{(t)}_{t+1}.$
The last step $\hat{w}^{(t)}_{t+1}=w^{(t)}_{t+1}$ follows from what we just proved in Step 1 that $\hat{w}^{(t)}_{\geq t+1}={w}^{(t)}_{\geq t+1}$. Thus we know
$\hat{q}_{t+1}=\mathcal{Q}(\hat{w}^{(t)}_{t+1})=\mathcal{Q}(w^{(t)}_{t+1})=q_{t+1}.$
This completes our induction. ∎